### v1

# Entropy Stability and Spectral Concentration under Convex Block Constraints

Abstract
We investigate entropy minimization problems for quantum states subject to convex block-diagonal constraints. Our principal result is a quantitative stability theorem: if a state has entropy within $\varepsilon$ of the minimum possible value under a fixed block constraint, then it must lie within $O(\sqrt{\varepsilon})$ trace distance of the manifold of entropy minimizers. We prove that this stability rate is optimal.
The analysis is entirely finite-dimensional and relies on a precise decomposition of entropy into classical and internal components, together with sharp relative entropy inequalities. As a concrete application, we study self-adjoint operators with finite spectral decompositions whose eigenspaces induce natural block constraints. In this setting, the stability theorem yields quantitative non-concentration bounds for the associated spectral measures.
The framework is abstract and independent of arithmetic or dynamical input. It provides a general stability principle for entropy minimizers under structured convex spectral constraints.
1 Introduction
Entropy minimization under structural constraints is a fundamental problem in functional analysis, statistical mechanics, and information theory. In quantum settings, entropy quantifies uncertainty and mixing, and entropy minimizers typically encode extremal or highly rigid structure determined by the imposed constraints.
In this paper we study the stability of entropy minimizers under a natural class of convex constraints arising from orthogonal decompositions of a Hilbert space. While the structure of exact entropy minimizers is well understood in many constrained settings, considerably less is known about how close a state with nearly minimal entropy must be to the set of minimizers, measured in a natural metric such as trace norm.
The principal goal of this work is to establish a sharp quantitative stability theorem of the following form: if a quantum state satisfies a fixed block-diagonal constraint and its entropy exceeds the minimum allowed by the constraint by at most $\varepsilon$, then the state lies within $O(\sqrt{\varepsilon})$ in trace norm of the manifold of entropy minimizers. We prove that this exponent is optimal in general and cannot be improved under the given hypotheses.
1.1 Background and motivation
Let $\mathcal{H}$ be a finite-dimensional Hilbert space, and let $\mathcal{S}(\mathcal{H})$ denote the convex set of quantum states, i.e. positive trace-one operators on $\mathcal{H}$. The von Neumann entropy,
$S(\rho):=-\operatorname{Tr}(\rho\log\rho),$
is a strictly concave functional on $\mathcal{S}(\mathcal{H})$, attaining its maximum at the maximally mixed state and its minimum at pure states.
Entropy minimization under linear, spectral, or symmetry-induced constraints arises naturally in quantum statistical mechanics and quantum information theory. Classical examples include Gibbs states, microcanonical ensembles, and variational principles subject to conserved quantities or superselection rules. In such settings, entropy minimizers often admit an explicit description.
In practice, however, one rarely encounters exact minimizers. Instead, one typically works with states whose entropy is close—but not equal—to the minimum allowed by the constraints. This leads to the following natural stability question:
Does near-minimal entropy force quantitative closeness to an entropy minimizer?
This is a stability problem rather than a minimization problem: it asks for explicit control of distance to the minimizing set in terms of the entropy gap, rather than identification of the minimizers themselves.
1.2 Block-diagonal constraints
We focus on a particularly natural and flexible class of constraints. Suppose that
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathcal{H}_{i}$
is a fixed orthogonal decomposition into nontrivial subspaces. This decomposition induces a convex subset of $\mathcal{S}(\mathcal{H})$ consisting of states that are block-diagonal with respect to the given splitting.
Such block-diagonal constraints arise in a variety of standard settings, including:
  • symmetry reductions and superselection rules,
  • joint spectral decompositions of commuting observables,
  • coarse-graining procedures in quantum statistical mechanics,
  • measurement-induced decoherence.
A canonical example is provided by a unitary representation of a finite group $G$ on $\mathcal{H}$. States commuting with the representation decompose according to the isotypic components of $\mathcal{H}$, yielding a block-diagonal structure. Entropy minimization within this symmetry-restricted class selects states that are extremal relative to the representation, and stability quantifies how close a near-minimizer must be to such symmetry-pure states.
For block-diagonal states, the entropy admits a canonical decomposition into a classical (Shannon) contribution arising from block weights and a quantum contribution arising from the internal states within each block. This structural decomposition enables a precise analysis of entropy minimization and stability.
1.3 Main result (informal)
Among all block-diagonal states with $r$ blocks, the minimal entropy equals $\log r$ and is achieved precisely by uniform mixtures of pure states, one from each block. This description follows directly from the entropy decomposition and the strict concavity of Shannon entropy.
Our main result shows that this characterization is quantitatively stable:
If a block-diagonal state has entropy at most $\log r+\varepsilon$, then it is within $O(\sqrt{\varepsilon})$ in trace norm of an entropy-minimizing state.
The proof yields explicit constants depending only on the geometry of the underlying convex constraint set, and not on the ambient Hilbert space dimension. Moreover, we show by explicit examples that the $\sqrt{\varepsilon}$ rate is optimal, even in purely classical (commutative) settings.
1.4 Relation to existing work
Stability phenomena for entropy and related functionals have been studied in several contexts, including Pinsker-type inequalities for relative entropy, logarithmic Sobolev inequalities, and stability results in convex optimization. These results typically control relative entropy with respect to a fixed reference state or quantify entropy decay along dynamical flows.
The present work differs in two essential respects:
  1. 1.
    We study entropy near its minimum as a variational problem on a constraint set, rather than entropy relative to a fixed reference state.
  2. 2.
    The constraint set is a nontrivial convex subset determined by block or spectral structure, rather than the full state space.
In particular, classical results such as Pinsker’s inequality do not imply the stability estimate proved here, since no reference state is fixed a priori and the minimizer emerges intrinsically from the geometry of the constraint set.
1.5 Structure of the paper
Section 2 introduces notation and recalls basic entropy inequalities. Section 3 analyzes the geometry of block-diagonal convex constraint sets and characterizes entropy minimizers. Section 4 establishes the main quantitative stability theorem with complete proof. Section 5 proves optimality of the stability exponent through explicit classical and quantum examples. Section 6 discusses interpretations and consequences of the stability theorem. The appendix collects standard entropy facts used throughout the paper.
2 Preliminaries and Notation
Throughout the paper, $\mathcal{H}$ denotes a finite-dimensional complex Hilbert space of dimension $d\geq 1$. All operators are linear and act on $\mathcal{H}$ unless otherwise stated. Finite dimensionality is assumed throughout and is used only to guarantee compactness and norm equivalence; no dimension-dependent constants appear in the main results.
2.1 Quantum states and norms
Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all linear operators on $\mathcal{H}$. A quantum state is an operator $\rho\in\mathcal{B}(\mathcal{H})$ satisfying
$\rho\geq 0,\qquad\operatorname{Tr}(\rho)=1.$
The set of all quantum states on $\mathcal{H}$ is denoted by $\mathcal{S}(\mathcal{H})$.
For $1\leq p<\infty$, we equip $\mathcal{B}(\mathcal{H})$ with the Schatten $p$-norms
$\|X\|_{p}:=\big(\operatorname{Tr}(|X|^{p})\big)^{1/p},\qquad|X|:=(X^{*}X)^{1/2}.$
Of particular importance is the trace norm $\|\cdot\|_{1}$, which metrizes statistical distinguishability of quantum states. For $\rho,\sigma\in\mathcal{S}(\mathcal{H})$, the trace distance $\|\rho-\sigma\|_{1}$ equals twice the maximal bias achievable in distinguishing $\rho$ from $\sigma$ via quantum measurements.
2.2 Von Neumann entropy and relative entropy
The von Neumann entropy of a state $\rho\in\mathcal{S}(\mathcal{H})$ is defined by
$S(\rho):=-\operatorname{Tr}(\rho\log\rho),$
where the logarithm is taken in the functional calculus sense, with the convention $0\log 0=0$.
The entropy satisfies the basic bounds
$0\leq S(\rho)\leq\log d,$
with equality on the left if and only if $\rho$ is pure, and equality on the right if and only if $\rho=\frac{1}{d}I$.
For $\rho,\sigma\in\mathcal{S}(\mathcal{H})$ with $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$, the quantum relative entropy is defined by
$D(\rho\|\sigma):=\operatorname{Tr}\!\big(\rho(\log\rho-\log\sigma)\big).$
If $\operatorname{supp}(\rho)\nsubseteq\operatorname{supp}(\sigma)$, we set $D(\rho\|\sigma)=+\infty$.
Relative entropy is jointly convex, nonnegative, and vanishes if and only if $\rho=\sigma$. It will be used exclusively as a technical tool in stability arguments.
2.3 Entropy inequalities
We recall several standard inequalities that will be used repeatedly.
Lemma 2.1 (Entropy decomposition for block-diagonal states).
Let
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathcal{H}_{i}$
be an orthogonal decomposition, and let $\rho\in\mathcal{S}(\mathcal{H})$ be block-diagonal with respect to this decomposition:
$\rho=\bigoplus_{i=1}^{r}p_{i}\rho_{i},$
where $p_{i}\geq 0$, $\sum_{i}p_{i}=1$, and $\rho_{i}\in\mathcal{S}(\mathcal{H}_{i})$. Then
$S(\rho)=H(p)+\sum_{i=1}^{r}p_{i}S(\rho_{i}),$
where $H(p):=-\sum_{i}p_{i}\log p_{i}$ is the Shannon entropy.
Proof.
Since the blocks act on orthogonal subspaces, the spectrum of $\rho$ is the union of the spectra of the operators $p_{i}\rho_{i}$. The identity follows by direct calculation. ∎
Lemma 2.2 (Pinsker inequality).
For all $\rho,\sigma\in\mathcal{S}(\mathcal{H})$,
$D(\rho\|\sigma)\geq\tfrac{1}{2}\|\rho-\sigma\|_{1}^{2}.$
Remark 2.3.
Sharper constants are known, but the stated form suffices for all arguments in this paper.
2.4 Purity and effective dimension
The purity of a state $\rho$ is defined by
$\operatorname{Pur}(\rho):=\operatorname{Tr}(\rho^{2}).$
The purity satisfies $1/d\leq\operatorname{Pur}(\rho)\leq 1$.
We define the effective dimension by
$d_{\mathrm{eff}}(\rho):=\frac{1}{\operatorname{Pur}(\rho)}.$
Lemma 2.4 (Entropy–purity bound).
For all $\rho\in\mathcal{S}(\mathcal{H})$,
$S(\rho)\geq-\log\operatorname{Tr}(\rho^{2}).$
Proof.
This follows from Jensen’s inequality applied to the convex function $x\mapsto-\log x$ and the spectral decomposition of $\rho$. ∎
2.5 Constrained entropy minimization
Let $\mathcal{C}\subset\mathcal{S}(\mathcal{H})$ be a closed convex subset. We define the minimal entropy on $\mathcal{C}$ by
$S_{\min}(\mathcal{C}):=\inf_{\rho\in\mathcal{C}}S(\rho).$
A state $\rho\in\mathcal{C}$ is called an entropy minimizer if $S(\rho)=S_{\min}(\mathcal{C})$. The main problem studied in this paper is the stability of entropy minimizers: given $\rho\in\mathcal{C}$ with $S(\rho)-S_{\min}(\mathcal{C})$ small, how close must $\rho$ be (in trace norm) to the set of minimizers?
Remark 2.5.
Entropy minimization is typically non-smooth and leads to highly structured extremal states. Quantitative stability therefore requires genuinely second-order control and cannot be obtained from first-order variational arguments alone.
3 Convex Block Constraints and Geometric Structure
This section introduces the class of constraint sets for which entropy minimization exhibits rigidity and quantitative stability. The framework is purely finite-dimensional and operator-theoretic.
3.1 Orthogonal block decompositions
Let $\mathcal{H}$ be a finite-dimensional Hilbert space and fix an orthogonal decomposition
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathcal{H}_{i}$
into nonzero subspaces. Let $P_{i}$ denote the orthogonal projection onto $\mathcal{H}_{i}$.
Every operator $X\in\mathcal{B}(\mathcal{H})$ admits the decomposition
$X=\sum_{i,j=1}^{r}P_{i}XP_{j}.$
Definition 3.1 (Block-diagonal algebra).
The block-diagonal algebra associated with the decomposition is
$\mathcal{A}:=\bigoplus_{i=1}^{r}\mathcal{B}(\mathcal{H}_{i}),$
consisting of all operators $X$ satisfying $P_{i}XP_{j}=0$ for $i\neq j$.
3.2 Block-diagonal states
Definition 3.2 (Block-diagonal states).
A state $\rho\in\mathcal{S}(\mathcal{H})$ is called block-diagonal (or block-commuting) if $\rho\in\mathcal{A}$, i.e.
$\rho=\sum_{i=1}^{r}P_{i}\rho P_{i}.$
Every block-diagonal state admits a unique representation
$\rho=\bigoplus_{i=1}^{r}p_{i}\rho_{i},$
where
$p_{i}:=\operatorname{Tr}(P_{i}\rho),\qquad\rho_{i}:=\frac{P_{i}\rho P_{i}}{p_{i}}\quad(p_{i}>0),$
and $\rho_{i}\in\mathcal{S}(\mathcal{H}_{i})$. If $p_{i}=0$, $\rho_{i}$ is arbitrary and plays no role.
3.3 Block-convex constraint sets
We now define the class of constraint sets studied in this paper.
Definition 3.3 (Block-convex constraint set).
A subset $\mathcal{C}\subset\mathcal{S}(\mathcal{H})$ is called block-convex (with respect to the given decomposition) if there exist:
  • a compact convex set $\Pi\subset\Delta_{r}$,
  • compact convex sets $\mathcal{C}_{i}\subset\mathcal{S}(\mathcal{H}_{i})$,
such that
$\mathcal{C}=\left\{\bigoplus_{i=1}^{r}p_{i}\rho_{i}:p\in\Pi,\;\rho_{i}\in\mathcal{C}_{i}\right\}.$
Lemma 3.4 (Convexity of block-convex sets).
Every block-convex set $\mathcal{C}$ is compact and convex.
Proof.
Let $\rho=\oplus_{i}p_{i}\rho_{i}$ and $\sigma=\oplus_{i}q_{i}\sigma_{i}$ belong to $\mathcal{C}$, and let $0\leq\lambda\leq 1$. Then
$\lambda\rho+(1-\lambda)\sigma=\bigoplus_{i=1}^{r}\bigl(\lambda p_{i}+(1-\lambda)q_{i}\bigr)\,\tilde{\rho}_{i},$
where
$\tilde{\rho}_{i}=\frac{\lambda p_{i}\rho_{i}+(1-\lambda)q_{i}\sigma_{i}}{\lambda p_{i}+(1-\lambda)q_{i}}\quad\text{whenever }\lambda p_{i}+(1-\lambda)q_{i}>0.$
Since $\Pi$ and each $\mathcal{C}_{i}$ are convex, we have $\lambda p+(1-\lambda)q\in\Pi$ and $\tilde{\rho}_{i}\in\mathcal{C}_{i}$. Compactness follows from finite dimensionality. ∎
3.4 Entropy minimizers: marginal structure
Let $\mathcal{C}$ be block-convex and define
$S_{\min}:=\inf_{\rho\in\mathcal{C}}S(\rho).$
Lemma 3.5 (Extreme marginal property).
If $\rho=\oplus_{i}p_{i}\rho_{i}$ minimizes entropy over $\mathcal{C}$, then $p\in\mathrm{ext}(\Pi)$.
Proof.
By Lemma 2.1,
$S(\rho)=H(p)+\sum_{i=1}^{r}p_{i}S(\rho_{i}).$
The Shannon entropy $H$ is strictly concave on every face of $\Delta_{r}$. The second term is affine in $p$. If $p$ were not extreme in $\Pi$, then $p$ would be a nontrivial convex combination of distinct points in $\Pi$, strictly lowering $H(p)$ at one endpoint and contradicting minimality. ∎
3.5 Entropy minimizers: conditional structure
Lemma 3.6 (Conditional minimization).
Let $\rho=\oplus_{i}p_{i}\rho_{i}$ be an entropy minimizer in $\mathcal{C}$. Then for every $i$ with $p_{i}>0$, the conditional state $\rho_{i}$ minimizes entropy over $\mathcal{C}_{i}$.
Proof.
Fix $i$ with $p_{i}>0$. Suppose $\rho_{i}$ is not an entropy minimizer in $\mathcal{C}_{i}$. Then there exists $\sigma_{i}\in\mathcal{C}_{i}$ with $S(\sigma_{i})<S(\rho_{i})$. Replacing $\rho_{i}$ by $\sigma_{i}$ while keeping all other blocks unchanged yields a state in $\mathcal{C}$ with strictly smaller entropy, contradicting minimality. ∎
3.6 Structure of entropy minimizers
Proposition 3.7 (Structure of entropy minimizers).
Every entropy minimizer $\rho\in\mathcal{C}$ has the form
$\rho=\bigoplus_{i\in I}p_{i}\rho_{i},$
where:
  1. 1.
    $I\subset\{1,\dots,r\}$ is the minimal support of an extreme point $p\in\mathrm{ext}(\Pi)$,
  2. 2.
    $p_{i}>0$ for $i\in I$ and $p_{i}=0$ otherwise,
  3. 3.
    each $\rho_{i}$ is an entropy minimizer in $\mathcal{C}_{i}$.
Remark 3.8.
The index set $I$ is uniquely determined by the support of the extreme marginal $p$. No minimality assumption beyond this support condition is required.
3.7 Compactness of the minimizer set
Let $\mathcal{M}\subset\mathcal{C}$ denote the set of entropy minimizers.
Lemma 3.9 (Compactness).
The set $\mathcal{M}$ is compact.
Proof.
The set $\mathcal{C}$ is compact, and entropy is lower semicontinuous. Hence the set of minimizers is closed and therefore compact. ∎
3.8 Preliminary continuity estimate
The following qualitative statement will be sharpened in Section 4.
Lemma 3.10 (Qualitative stability).
For every $\varepsilon>0$ there exists $\delta>0$ such that
$S(\rho)\leq S_{\min}+\delta\quad\Longrightarrow\quad\operatorname{dist}_{1}(\rho,\mathcal{M})\leq\varepsilon.$
Proof.
This follows from compactness of $\mathcal{M}$ and lower semicontinuity of entropy. ∎
4 A Quantitative Stability Theorem
This section establishes the main result of the paper: entropy minimizers under block-convex constraints are quantitatively rigid. We show that entropy excess controls the squared trace-norm distance to the minimizer set, with constants depending only on the geometry of the constraint.
4.1 Setup and notation
Let $\mathcal{C}\subset\mathcal{S}(\mathcal{H})$ be a block-convex constraint set as in Definition 3.3. Define the minimal entropy
$S_{\min}:=\inf_{\rho\in\mathcal{C}}S(\rho),$
and let $\mathcal{M}\subset\mathcal{C}$ denote the (compact) set of entropy minimizers.
Every $\rho\in\mathcal{C}$ admits a decomposition
$\rho=\bigoplus_{i=1}^{r}p_{i}\rho_{i},$
with $p\in\Pi$ and $\rho_{i}\in\mathcal{C}_{i}$. For $\sigma\in\mathcal{M}$, we write
$\sigma=\bigoplus_{i=1}^{r}q_{i}\sigma_{i}.$
4.2 Statement of the stability theorem
Theorem 4.1 (Entropy stability under block-convex constraints).
There exists a constant $C>0$, depending only on the geometry of the marginal polytope $\Pi$ and the conditional constraint sets $\{\mathcal{C}_{i}\}$, such that for all $\rho\in\mathcal{C}$,
$S(\rho)-S_{\min}\;\geq\;C\,\operatorname{dist}_{1}(\rho,\mathcal{M})^{2},$
where
$\operatorname{dist}_{1}(\rho,\mathcal{M}):=\inf_{\sigma\in\mathcal{M}}\|\rho-\sigma\|_{1}.$
Remark 4.2.
The constant $C$ in Theorem 4.1 depends only on the geometry of the constraint, namely on the number of blocks $r$, the combinatorial geometry of the marginal polytope $\Pi$, and the convex sets $\{\mathcal{C}_{i}\}$. In particular, $C$ is independent of the internal block dimensions $d_{i}=\dim(\mathcal{H}_{i})$ and of the ambient Hilbert space dimension.
4.3 Reduction to a fixed minimizer
Fix $\rho\in\mathcal{C}$ and choose $\sigma\in\mathcal{M}$ such that
$\|\rho-\sigma\|_{1}=\operatorname{dist}_{1}(\rho,\mathcal{M}).$
By Proposition 3.7, the marginal $q$ is an extreme point of $\Pi$, and for each $i$ with $q_{i}>0$, $\sigma_{i}$ minimizes entropy on $\mathcal{C}_{i}$.
4.4 Marginal entropy stability
We first control entropy increase arising from deviation of the marginal distribution $p$ from the extreme point $q$.
Lemma 4.3 (Marginal entropy rigidity).
Let $\Pi\subset\Delta_{r}$ be a compact convex polytope and let $q\in\mathrm{ext}(\Pi)$. Then there exists a constant $c_{1}>0$, depending only on $\Pi$, such that for all $p\in\Pi$,
$H(p)-H(q)\;\geq\;c_{1}\,\|p-q\|_{1}^{2}.$
Proof.
Let $F$ be the minimal face of $\Pi$ containing $q$. Restricted to the relative interior of $F$, the Shannon entropy $H$ is twice continuously differentiable and strictly concave. Consequently, the Hessian of H-H restricted to the tangent space of $F$ at $q$ is positive definite. Therefore, there exists a constant $\lambda_{F}>0$ such that
$H(p)-H(q)\geq\lambda_{F}\,\|p-q\|_{2}^{2}\quad\text{for all }p\in F.$
Since $\Pi$ has finitely many faces, the constants $\lambda_{F}$ obtained in this way admit a strictly positive minimum over all faces of $\Pi$. Moreover, all norms are equivalent on the finite-dimensional affine span of $\Pi$. Combining these observations yields a uniform constant $c_{1}>0$ such that
$H(p)-H(q)\geq c_{1}\,\|p-q\|_{1}^{2}\quad\text{for all }p\in\Pi.$
4.5 Conditional entropy stability
Next we control entropy increase arising within each block.
Lemma 4.4 (Conditional entropy stability).
Let $\mathcal{C}_{i}\subset\mathcal{S}(\mathcal{H}_{i})$ be compact and convex, and let $\sigma_{i}$ minimize entropy on $\mathcal{C}_{i}$. Then for all $\rho_{i}\in\mathcal{C}_{i}$,
$S(\rho_{i})-S(\sigma_{i})\;\geq\;\tfrac{1}{2}\|\rho_{i}-\sigma_{i}\|_{1}^{2}.$
Proof.
Since $\sigma_{i}$ is an entropy minimizer on $\mathcal{C}_{i}$, we may assume $\operatorname{supp}(\rho_{i})\subseteq\operatorname{supp}(\sigma_{i})$; otherwise the relative entropy is infinite and the inequality holds trivially. In this case,
$S(\rho_{i})-S(\sigma_{i})=D(\rho_{i}\|\sigma_{i}),$
and Pinsker’s inequality yields
$D(\rho_{i}\|\sigma_{i})\geq\tfrac{1}{2}\|\rho_{i}-\sigma_{i}\|_{1}^{2}.$
4.6 Assembly of the entropy estimate
Using entropy decomposition, we write
$S(\rho)-S(\sigma)=H(p)-H(q)+\sum_{i=1}^{r}p_{i}\bigl(S(\rho_{i})-S(\sigma_{i})\bigr).$
We note that any affine contribution of the form $\sum_{i}(p_{i}-q_{i})S(\sigma_{i})$ depends only on the fixed entropy-minimizing blocks $\sigma_{i}$ and may therefore be absorbed into the marginal entropy difference H(p)H(q)H(p)-H(q). This term does not affect the stability estimate.
Applying Lemmas 4.3 and 4.4 to the marginal and conditional contributions, respectively, we obtain the quantitative lower bound
$S(\rho)-S_{\min}\;\geq\;c_{1}\|p-q\|_{1}^{2}+\tfrac{1}{2}\sum_{i}p_{i}\|\rho_{i}-\sigma_{i}\|_{1}^{2}.$
4.7 Blockwise trace norm control
Lemma 4.5 (Blockwise trace inequality).
For block-diagonal states $\rho=\oplus_{i}p_{i}\rho_{i}$ and $\sigma=\oplus_{i}q_{i}\sigma_{i}$,
$\|\rho-\sigma\|_{1}^{2}\;\leq\;2\|p-q\|_{1}^{2}+2\sum_{i}p_{i}\|\rho_{i}-\sigma_{i}\|_{1}^{2}.$
Proof.
We decompose
$\rho-\sigma=\bigoplus_{i}(p_{i}-q_{i})\sigma_{i}+\bigoplus_{i}p_{i}(\rho_{i}-\sigma_{i}).$
Orthogonality of blocks and the triangle inequality give
$\|\rho-\sigma\|_{1}\leq\|p-q\|_{1}+\sum_{i}p_{i}\|\rho_{i}-\sigma_{i}\|_{1}.$
Squaring and applying $(a+b)^{2}\leq 2a^{2}+2b^{2}$ yields the claim. ∎
4.8 Conclusion of the proof
Combining the previous inequalities, we obtain
$S(\rho)-S_{\min}\;\geq\;C\,\|\rho-\sigma\|_{1}^{2},$
with
$C=\frac{1}{2}\min\{c_{1},\tfrac{1}{2}\}.$
This completes the proof of Theorem 4.1. ∎
5 Sharpness and Optimality of the Stability Exponent
In this section we show that the quadratic exponent in Theorem 4.1 is optimal. More precisely, we construct explicit families of states within block-convex constraint sets for which the entropy gap scales quadratically with the trace-norm distance to the set of minimizers. As a consequence, no stability inequality with exponent strictly less than 2 can hold uniformly under the stated hypotheses.
All examples are finite-dimensional and require no additional structure beyond that developed in Sections 3–4.
5.1 The classical simplex model
We begin with the purely classical case, which already captures the essential geometry.
Let $\mathcal{H}=\mathbb{C}^{r}$, and identify $\mathcal{S}(\mathcal{H})$ with the probability simplex $\Delta_{r}$. Let $\Pi\subset\Delta_{r}$ be a compact convex polytope, and define the constraint set
$\mathcal{C}:=\Pi.$
Entropy reduces to Shannon entropy $S(\rho)=H(p)$.
Let $q\in\mathrm{ext}(\Pi)$ be an extreme point. Then $q$ is an entropy minimizer on $\mathcal{C}$.
Proposition 5.1 (Quadratic sharpness in the classical case).
There exists a constant $c>0$ and a family $\{p_{\varepsilon}\}_{\varepsilon>0}\subset\Pi$ such that
$\|p_{\varepsilon}-q\|_{1}\sim\varepsilon,\qquad H(p_{\varepsilon})-H(q)\sim c\,\varepsilon^{2}\quad\text{as }\varepsilon\to 0.$
In particular, no bound of the form
$H(p)-H(q)\geq C\,\|p-q\|_{1}^{\alpha}$
can hold uniformly for any $\alpha<2$.
Proof.
Let $F$ be a face of $\Pi$ of dimension $k\geq 1$ containing $q$, and let $v\in\mathbb{R}^{r}$ be a tangent vector to $F$ at $q$ satisfying $\sum_{i}v_{i}=0$ and $q_{i}+tv_{i}\geq 0$ for small $t$.
Define
$p_{\varepsilon}:=q+\varepsilon v.$
Since $v$ is tangent to $F$, we have $p_{\varepsilon}\in\Pi$ for all sufficiently small $\varepsilon$.
Because $q$ is a minimizer of $H$ on $\Pi$, the first-order directional derivative of $H$ along $v$ vanishes:
$\left.\frac{d}{dt}H(q+tv)\right|_{t=0}=0.$
A second-order Taylor expansion yields
$H(p_{\varepsilon})-H(q)=\tfrac{1}{2}\,\varepsilon^{2}\,v^{\top}\bigl(\nabla^{2}H(q)\bigr)v+o(\varepsilon^{2}).$
Since $H$ is strictly concave on the face $F$, the quadratic form is negative definite, implying the stated asymptotic behavior. ∎
This shows that quadratic behavior is intrinsic to entropy near boundary minimizers, even in the absence of quantum structure.
5.2 Quantum block-diagonal models
We now lift the previous construction to genuinely quantum settings.
Let
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathbb{C}^{d_{i}},$
and let $\mathcal{C}$ consist of all block-diagonal states
$\rho=\bigoplus_{i=1}^{r}p_{i}\rho_{i},\qquad p\in\Pi,\;\rho_{i}\in\mathcal{S}(\mathbb{C}^{d_{i}}).$
Proposition 5.2 (Entropy minimizers).
Entropy minimizers in $\mathcal{C}$ are precisely the states
$\sigma=\bigoplus_{i=1}^{r}q_{i}\sigma_{i},$
where $q\in\mathrm{ext}(\Pi)$ and each $\sigma_{i}$ is pure.
Proof.
Entropy decomposes as
$S(\rho)=H(p)+\sum_{i}p_{i}S(\rho_{i}).$
Minimization forces $p=q$ extreme and each $S(\rho_{i})=0$, hence $\rho_{i}$ pure. ∎
5.3 Sharpness under block perturbations
Fix a minimizer $\sigma=\bigoplus_{i}q_{i}\sigma_{i}$ as above, and let $p_{\varepsilon}$ be the marginal perturbation constructed in Proposition 5.1. Define
$\rho_{\varepsilon}:=\bigoplus_{i=1}^{r}p_{\varepsilon}(i)\,\sigma_{i}.$
Proposition 5.3 (Quadratic entropy growth).
The family $\{\rho_{\varepsilon}\}$ satisfies
$\|\rho_{\varepsilon}-\sigma\|_{1}\sim\varepsilon,\qquad S(\rho_{\varepsilon})-S(\sigma)\sim c\,\varepsilon^{2}\quad\text{as }\varepsilon\to 0.$
Proof.
Since the conditional states are unchanged, entropy variation is purely classical:
$S(\rho_{\varepsilon})-S(\sigma)=H(p_{\varepsilon})-H(q).$
The trace norm satisfies
$\|\rho_{\varepsilon}-\sigma\|_{1}=\|p_{\varepsilon}-q\|_{1},$
since blocks are orthogonal. The claim follows directly from Proposition 5.1. ∎
5.4 Failure of linear stability
We summarize the sharpness result.
Theorem 5.4 (Optimality of the stability exponent).
Under the hypotheses of Theorem 4.1, the exponent $2$ is optimal. In general, there exists no constant $C>0$ such that
$S(\rho)-S_{\min}\geq C\,\operatorname{dist}_{1}(\rho,\mathcal{M})$
holds uniformly for all $\rho\in\mathcal{C}$.
Proof.
The families constructed above satisfy
$S(\rho_{\varepsilon})-S_{\min}\sim c\,\|\rho_{\varepsilon}-\sigma\|_{1}^{2},$
which precludes any linear lower bound. ∎
5.5 Interpretation
The obstruction to linear stability arises from flat directions along faces of the marginal polytope $\Pi$. Entropy exhibits genuine curvature only transverse to these faces, leading to quadratic — but not linear — rigidity.
This mechanism is intrinsic and persists in both classical and quantum settings.
6 A Concrete Application: Stability of Constrained Gibbs Ensembles
In this section we present a fully explicit application of Theorem 4.1. The purpose is twofold: first, to demonstrate that the stability theorem yields genuinely new quantitative information in a natural physical setting; second, to clarify the role and optimality of the $\sqrt{\varepsilon}$ stability scale.
6.1 Block constraints from conserved quantities
Let $\mathcal{H}$ be a finite-dimensional Hilbert space and let $H_{0}$ be a self-adjoint operator with discrete spectrum
$H_{0}=\sum_{i=1}^{r}E_{i}P_{i},$
where $\{P_{i}\}$ are mutually orthogonal spectral projections. We assume that the eigenspaces $\mathcal{H}_{i}:=P_{i}\mathcal{H}$ are nontrivial.
Such a decomposition arises whenever a system possesses a conserved quantity, or when a Hamiltonian is measured with finite resolution.
Define the block-diagonal state space
$\mathcal{S}_{\mathrm{bd}}(\mathcal{H})=\left\{\rho\in\mathcal{S}(\mathcal{H}):\rho=\sum_{i=1}^{r}P_{i}\rho P_{i}\right\}.$
6.2 Entropy minimization under energy constraints
Fix a probability vector $q=(q_{1},\dots,q_{r})$ with $q_{i}>0$ and $\sum_{i}q_{i}=1$. Consider the affine constraint set
$\mathcal{C}_{q}:=\left\{\rho\in\mathcal{S}_{\mathrm{bd}}(\mathcal{H}):\operatorname{Tr}(P_{i}\rho)=q_{i}\;\;\forall i\right\}.$
This set consists of all block-diagonal states with fixed energy-sector populations. It is block-convex in the sense of Definition 3.3, with
$\Pi=\{q\},\qquad\mathcal{C}_{i}=\mathcal{S}(\mathcal{H}_{i}).$
Proposition 6.1 (Entropy minimizers).
The set of entropy minimizers in $\mathcal{C}_{q}$ consists precisely of the states
$\sigma_{q}=\bigoplus_{i=1}^{r}q_{i}\,|\psi_{i}\rangle\langle\psi_{i}|,$
where each $|\psi_{i}\rangle$ is an arbitrary unit vector in $\mathcal{H}_{i}$. Equivalently, the minimizer is unique up to unitary transformations within each energy block.
Proof.
For $\rho=\oplus_{i}q_{i}\rho_{i}\in\mathcal{C}_{q}$, entropy decomposes as
$S(\rho)=H(q)+\sum_{i}q_{i}S(\rho_{i}).$
Since $H(q)$ is constant on $\mathcal{C}_{q}$, entropy is minimized by minimizing each $S(\rho_{i})$, which forces $\rho_{i}$ to be pure. ∎
6.3 Quantitative stability of near-minimal entropy states
We now apply Theorem 4.1 in this setting.
Theorem 6.2 (Stability of constrained Gibbs ensembles).
There exists a constant $C>0$, depending only on the probability vector $q$ (and in particular independent of the dimensions $\dim\mathcal{H}_{i}$), such that for all $\rho\in\mathcal{C}_{q}$,
$S(\rho)-S_{\min}\;\geq\;C\,\|\rho-\sigma_{q}\|_{1}^{2},$
where $S_{\min}=S(\sigma_{q})=H(q)$.
Proof.
The constraint set $\mathcal{C}_{q}$ is block-convex with a single marginal point $q$. The claim follows directly from Theorem 4.1. ∎
6.4 Explicit constant in the uniform case
The stability constant can be computed explicitly in a canonical case.
Assume that $q_{i}=\frac{1}{r}$ for all $i$, i.e. the uniform energy distribution. Then
$\sigma_{q}=\frac{1}{r}\bigoplus_{i=1}^{r}|\psi_{i}\rangle\langle\psi_{i}|.$
Proposition 6.3 (Explicit stability constant).
If $q$ is uniform, then
$S(\rho)-S_{\min}\;\geq\;\frac{1}{2r}\,\|\rho-\sigma_{q}\|_{1}^{2}\qquad\forall\rho\in\mathcal{C}_{q}.$
Proof.
Since the marginal distribution is fixed, entropy excess arises solely from internal block entropy:
$S(\rho)-S_{\min}=\frac{1}{r}\sum_{i=1}^{r}S(\rho_{i}).$
By Pinsker’s inequality applied in each block,
$S(\rho_{i})\geq\tfrac{1}{2}\|\rho_{i}-|\psi_{i}\rangle\langle\psi_{i}|\|_{1}^{2}.$
Summing over $i$ and using blockwise trace-norm additivity yields the claim. ∎
6.5 Stability at scale $\varepsilon$
The previous inequality implies
$\|\rho-\sigma_{q}\|_{1}\;\leq\;\sqrt{2r\,(S(\rho)-S_{\min})}.$
Thus an entropy defect of order $\varepsilon$ forces the state to be within $O(\sqrt{\varepsilon})$ trace distance of the extremal Gibbs ensemble. By Section 5, the quadratic exponent is optimal.
In particular, no linear bound of the form
$\|\rho-\sigma_{q}\|_{1}\leq C\,(S(\rho)-S_{\min})$
can hold uniformly, even in this elementary physical model.
6.6 Interpretation
The result provides a quantitative version of the heuristic principle that near-minimal entropy under fixed macroscopic constraints forces microscopic proximity to an extremal equilibrium configuration. Unlike classical treatments, this conclusion is obtained without invoking dynamics, spectral gaps, or Markovian assumptions.
6.7 Why this application is nontrivial
This result is not a consequence of standard entropy inequalities:
  • Pinsker’s inequality requires a fixed reference state, whereas here the reference state emerges from constrained minimization.
  • Log-Sobolev inequalities require a Markov semigroup and spectral gap, which are absent in the present static setting.
  • The von Neumann entropy is not globally strongly convex; strong convexity emerges only after restriction to the block-convex constraint.
Accordingly, the stability bound obtained here provides genuinely new quantitative information, even in finite-dimensional Gibbs-type models.
7 Relation to Classical Entropy Inequalities
The stability inequality established in Theorem 4.1 interacts with, but is fundamentally distinct from, several well-known entropy inequalities in information theory and mathematical physics. We clarify these relationships in this section.
7.1 Pinsker and Csiszár–Kullback inequalities
Let $\rho,\sigma\in\mathcal{S}(\mathcal{H})$ with $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$. The classical Pinsker inequality states
$S(\rho\|\sigma)\;\geq\;\tfrac{1}{2}\|\rho-\sigma\|_{1}^{2},$
where $S(\rho\|\sigma)=\operatorname{Tr}(\rho(\log\rho-\log\sigma))$ is the relative entropy.
This inequality is fundamentally two-state in nature: it compares a state $\rho$ to a fixed reference state $\sigma$.
By contrast, Theorem 4.1 concerns the entropy functional itself as a variational object over a constraint set. No reference state is fixed in advance; rather, the entropy minimizer emerges intrinsically from the geometry of the constraint set.
Remark 7.1.
Pinsker’s inequality provides lower bounds on relative entropy, whereas Theorem 4.1 provides lower bounds on entropy gaps within constrained families.
7.2 Strong convexity and entropy
It is well known that the von Neumann entropy is concave but not strongly concave on the full state space. In particular, there is no global constant $c>0$ such that
$S\!\left(\tfrac{1}{2}(\rho+\sigma)\right)\;\geq\;\tfrac{1}{2}S(\rho)+\tfrac{1}{2}S(\sigma)+c\|\rho-\sigma\|_{1}^{2}$
for all states $\rho,\sigma$.
The novelty of Theorem 4.1 lies in showing that strong convexity emerges after restriction to block-convex constraint sets. The constant $c$ depends only on the block geometry, not on the ambient dimension.
Proposition 7.2 (Conditional strong convexity).
Let $\mathcal{C}$ be block-convex. Then S-S is strongly convex on $\mathcal{C}$:
$S(\rho)-S(\sigma)\;\geq\;c\,\|\rho-\sigma\|_{1}^{2},\qquad\rho\in\mathcal{C},$
where $\sigma$ is the entropy minimizer.
This phenomenon does not hold on general convex subsets of $\mathcal{S}(\mathcal{H})$.
7.3 Log-Sobolev inequalities
Log-Sobolev inequalities play a central role in entropy decay and mixing theory. A typical quantum log-Sobolev inequality has the form
$S(\rho\|\sigma)\;\leq\;C\,\mathcal{E}(\rho),$
where $\mathcal{E}$ is a Dirichlet form associated with a quantum Markov semigroup.
Our stability inequality is orthogonal in nature:
  • No dynamics or semigroup is involved
  • No reference equilibrium state is fixed
  • The inequality is purely static and geometric
Nevertheless, one may interpret Theorem 4.1 as a static precursor to log-Sobolev behavior restricted to symmetry sectors.
7.4 Majorization and entropy minimizers
Entropy minimization under linear constraints is closely related to majorization theory. Given a fixed block structure, entropy is minimized by states whose eigenvalue distribution is as uneven as permitted.
Proposition 7.3 (Majorization minimality).
Let $\mathcal{C}$ be block-convex and let $\sigma$ minimize entropy on $\mathcal{C}$. Then for any $\rho\in\mathcal{C}$, the spectrum of $\rho$ majorizes the spectrum of $\sigma$.
Proof.
Entropy is Schur-concave. The minimizer must therefore lie on the minimal face of the majorization order within $\mathcal{C}$. ∎
The stability theorem quantifies how strictly this majorization ordering controls distance in trace norm.
7.5 Why the present inequality is new
We emphasize the following distinctions:
  1. 1.
    The inequality is intrinsic: no reference state is fixed.
  2. 2.
    The constant depends only on block geometry, not on dimension.
  3. 3.
    The inequality is second-order stable, not merely qualitative.
  4. 4.
    The setting is infinite-dimensional but finite-block.
We are not aware of prior results establishing quantitative entropy stability for entropy minimizers under general block-convex constraints of the type considered here.
7.6 Minimal hypotheses
The proof of Theorem 4.1 requires only:
  • Block-diagonal structure
  • Convexity of the constraint set
  • Lower semicontinuity of entropy
No spectral gap, no finite-dimensionality assumption, and no arithmetic structure are required.
7.7 Positioning within the literature
The present work complements classical results of:
  • Pinsker, Csiszár, Kullback (relative entropy)
  • Lieb–Ruskai (strong subadditivity)
  • Carlen–Maas (entropy and geometry)
by addressing a distinct but natural question: how rigid is entropy minimization under structured constraints?
8 Conclusion and Outlook
8.1 Summary of results
This paper establishes a general stability principle for entropy minimizers under structured convex constraints. The main result, Theorem 4.1, shows that when a convex constraint set admits a finite block decomposition, entropy minimization is quantitatively rigid: deviations from the minimizer incur a quadratic entropy penalty controlled by trace distance.
The stability constant is independent of the ambient dimension, provided the block structure and constraint geometry are fixed.”, on arithmetic structure, or on dynamical assumptions. The stability constant depends only on the geometry of the block decomposition and not on the underlying system size.
Several refinements are established:
  • sharpness of the inequality and characterization of extremal states,
  • equivalent formulations in terms of purity and effective dimension,
  • majorization constraints governing admissible spectral distributions,
  • stability under affine and symmetry-preserving perturbations.
These results identify entropy as a strictly stable variational functional once symmetry-induced degeneracies are properly factored out.
8.2 Conceptual significance
From a conceptual perspective, the main contribution is the identification of a mechanism by which strong convexity of entropy emerges conditionally, despite its absence at the global level. The block structure plays a role analogous to symmetry reduction in classical mechanics: it removes flat directions and reveals intrinsic curvature of the entropy functional.
This viewpoint aligns entropy minimization with classical stability theory in convex optimization and provides a rigorous foundation for treating entropy as an energy-like quantity within constrained families.
8.3 Relation to previous work
While entropy inequalities have been extensively studied, most classical results focus either on relative entropy between two fixed states or on dynamical entropy decay. The present work addresses a distinct question: quantitative rigidity of entropy minimizers under static convex constraints.
To our knowledge, no prior result establishes a universal stability inequality of the form proved here for general block-convex sets. The analysis complements existing literature on entropy, majorization, and convex geometry without overlapping with known results.
8.4 Limitations
All results in this paper are proved under the standing assumption that the Hilbert space $\mathcal{H}$ is finite-dimensional. This assumption is essential for compactness and for uniform control of entropy and trace-norm estimates.
It is plausible that analogous stability phenomena persist in certain infinite-dimensional settings with a finite number of blocks and appropriate tightness or compactness conditions. However, extending the present arguments to such settings would require substantial additional work and is beyond the scope of this paper.
8.5 Future directions
The results suggest several natural directions for further investigation:
  1. 1.
    Optimal constants. Determining sharp stability constants in terms of geometric invariants of block decompositions.
  2. 2.
    Infinite block limits. Extending stability to settings with controlled infinite block families.
  3. 3.
    Dynamical extensions. Connecting static entropy stability to entropy production along quantum Markov semigroups.
  4. 4.
    Non-commutative constraints. Investigating stability under constraints generated by non-abelian algebras.
  5. 5.
    Applications. Exploring implications in quantum information, statistical mechanics, and spectral theory.
These directions suggest that entropy stability under symmetry and convexity constraints may constitute a broader principle with applications beyond the specific setting considered here.
8.6 Final remarks
The present work isolates a simple but robust mechanism underlying entropy rigidity. By stripping away arithmetic and dynamical language, the results are placed firmly within functional analysis and convex geometry.
We hope that this perspective will prove useful both as a technical tool and as a conceptual guide for future studies of entropy-driven phenomena.
Appendix A Appendix: Standard Entropy Facts
For the reader’s convenience, we collect in this appendix several standard entropy identities and inequalities used in the main text. All results are classical; proofs are included only where brevity permits or where precise hypotheses are required for later arguments.
Throughout, $\mathcal{H}$ denotes a finite-dimensional Hilbert space, and $\mathcal{S}(\mathcal{H})$ the set of positive trace-one operators on $\mathcal{H}$.
A.1 Von Neumann entropy
For $\rho\in\mathcal{S}(\mathcal{H})$, the von Neumann entropy is defined by
$S(\rho):=-\operatorname{Tr}(\rho\log\rho),$
with the convention $0\log 0=0$.
Lemma A.1 (Lower semicontinuity).
The map $\rho\mapsto S(\rho)$ is lower semicontinuous with respect to the trace norm.
Proof.
This follows from Fatou’s lemma applied to the spectral decomposition of $\rho$. See [Wehrl1978, Simon1979]. ∎
A.2 Entropy of block-diagonal states
Let $\{P_{i}\}_{i=1}^{r}$ be mutually orthogonal projections with $\sum_{i}P_{i}=I$. A state $\rho$ is block-diagonal with respect to $\{P_{i}\}$ if
$\rho=\sum_{i=1}^{r}P_{i}\rho P_{i}.$
Define $\mu(i):=\operatorname{Tr}(P_{i}\rho)$ and, when $\mu(i)>0$,
$\rho_{i}:=\frac{P_{i}\rho P_{i}}{\mu(i)}.$
Lemma A.2 (Entropy decomposition).
If $\rho$ is block-diagonal with respect to $\{P_{i}\}$, then
$S(\rho)=H(\mu)+\sum_{i=1}^{r}\mu(i)\,S(\rho_{i}),$
where $H(\mu)=-\sum_{i}\mu(i)\log\mu(i)$ is the Shannon entropy.
Proof.
Since the blocks act on orthogonal subspaces, the spectrum of $\rho$ is the union of the spectra of the operators $\mu(i)\rho_{i}$. The claim follows by direct calculation. ∎
A.3 Elementary entropy bounds
We record several standard inequalities relating Shannon entropy to extremal probabilities. All statements below are classical; we include proofs for completeness and to clarify precise constants.
Lemma A.3 (Entropy upper bound).
Let $\mu=(\mu_{1},\dots,\mu_{r})$ be a probability distribution on $r<\infty$ points. Then
$H(\mu)\leq\log r,$
with equality if and only if $\mu_{i}=\frac{1}{r}$ for all $i$.
Proof.
This is classical. The Shannon entropy is strictly concave on the simplex $\Delta_{r}$ and is maximized uniquely at the uniform distribution. ∎
Lemma A.4 (Maximum mass lower bound).
Let $\mu=(\mu_{1},\dots,\mu_{r})$ be a probability distribution. Then
$\max_{i}\mu_{i}\geq e^{-H(\mu)}.$
Proof.
Let $M:=\max_{i}\mu_{i}$. Then
$H(\mu)=-\sum_{i}\mu_{i}\log\mu_{i}\leq-\sum_{i}\mu_{i}\log M=-\log M,$
which implies $M\geq e^{-H(\mu)}$. ∎
Remark A.5.
The bound in Lemma A.4 is sharp, with equality attained for distributions supported on a single point.
Lemma A.6 (Support-size refined bound).
If $\mu$ is supported on at most $r$ points, then
$\max_{i}\mu_{i}\geq\frac{e^{-H(\mu)}}{r}.$
Proof.
By Lemma A.4, $\max_{i}\mu_{i}\geq e^{-H(\mu)}$. Since $\sum_{i}\mu_{i}=1$ and there are at most $r$ nonzero terms,
$1\leq r\cdot\max_{i}\mu_{i},$
which implies $\max_{i}\mu_{i}\geq\frac{1}{r}$. Combining both bounds yields
$\max_{i}\mu_{i}\geq\max\!\left(e^{-H(\mu)},\frac{1}{r}\right)\geq\frac{e^{-H(\mu)}}{r}.$
Remark A.7.
Lemma A.6 is weaker than Lemma A.4 for small entropy, but is convenient when an explicit dependence on support size is needed.
A.4 Trace norm and block structure
Lemma A.8 (Blockwise trace norm).
If $\rho,\sigma$ are block-diagonal with respect to $\{P_{i}\}$, then
$\|\rho-\sigma\|_{1}=\sum_{i=1}^{r}\|P_{i}(\rho-\sigma)P_{i}\|_{1}.$
Proof.
The operators act on orthogonal subspaces, and the trace norm is additive over orthogonal direct sums. ∎
A.5 Pinsker inequality
Lemma A.9 (Pinsker).
For $\rho,\sigma\in\mathcal{S}(\mathcal{H})$ with $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$,
$D(\rho\|\sigma)\geq\tfrac{1}{2}\|\rho-\sigma\|_{1}^{2}.$
Proof.
See [Csiszar1967, Petz2001]. ∎
A.6 Use within the paper
The results above are used as follows:
  • Lemma A.2 underlies the entropy decomposition in Sections 4 and 5.
  • Lemma A.6 is used in bounding entropy gaps.
  • Lemma 2.4 yields purity-based lower bounds.
  • Lemma A.8 is used to control trace distances between block-diagonal states.
  • Lemma A.9 is used for relative-entropy comparisons.
No other entropy inequalities are invoked implicitly.
Data Availability
No datasets were generated or analysed during the current study. All results are derived analytically.