### v2

# Quadratic Stability of Entropy Minimizers under Block-Separable Convex Constraints

Abstract
We study quadratic stability of entropy minimization under convex constraints with block-separable structure. For constraint sets of the form $\mathcal{C}=\{\bigoplus_{i}p_{i}\rho_{i}:p\in\Pi,\ \rho_{i}\in\mathcal{C}_{i}\}$, we establish explicit stability estimates showing that, under a confining (fixed-support) hypothesis, the entropy gap controls the squared trace-norm distance to the set of entropy minimizers, with constants determined by the geometry of the constraint.
The stability constant admits a natural decomposition into marginal and conditional components: the marginal contribution is governed by the curvature of Shannon entropy on the marginal polytope $\Pi$ at extreme points, while the conditional contribution is governed by the curvature of von Neumann entropy at entropy-minimizing states within each block.
We prove that the quadratic exponent is optimal by constructing explicit examples for which no linear stability bound holds uniformly. Exploiting the block-separable structure, the analysis reduces the global stability problem to independent marginal and conditional subproblems, yielding a geometric characterization of the stability constants in terms of constraint curvature. This stability phenomenon cannot be derived from Pinsker–type inequalities or standard entropy continuity bounds, since no reference state is fixed and the minimizer emerges intrinsically from the constraint geometry.
keywords:
convex optimization, entropy minimization, quadratic stability, block-separable constraints, error bounds, explicit constants
{AMS}
90C25, 94A17, 49J52, 49J45
1 Introduction
Entropy minimization under structural constraints is a fundamental problem in functional analysis, statistical mechanics, and information theory. In quantum settings, entropy quantifies uncertainty and mixing, and entropy minimizers typically encode extremal or highly rigid structure determined by the imposed constraints.
We study the stability of entropy minimizers under a natural class of convex constraints arising from orthogonal decompositions of a Hilbert space. While the structure of exact entropy minimizers is well understood in many constrained settings, considerably less is known about how close a state with nearly minimal entropy must be to the set of minimizers, measured in a natural metric such as the trace norm.
The principal goal of this work is to establish a sharp quantitative stability theorem of the following form: if a quantum state satisfies a fixed block-diagonal constraint and its entropy exceeds the minimum allowed by the constraint by at most $\varepsilon$, then the state lies within $O(\sqrt{\varepsilon})$ in trace norm of the set of entropy minimizers. We prove that this exponent is optimal in general and cannot be improved under the given hypotheses. Purely quadratic entropy–distance stability holds under fixed-support (confining) constraints; see Appendix A for a precise geometric formulation.
1.1 Background and motivation
Let $\mathcal{H}$ be a finite-dimensional Hilbert space, and let $\mathcal{S}(\mathcal{H})$ denote the convex set of quantum states, i.e. positive trace-one operators on $\mathcal{H}$. The von Neumann entropy,
$S(\rho):=-\operatorname{Tr}(\rho\log\rho),$
is a strictly concave functional on $\mathcal{S}(\mathcal{H})$, attaining its maximum at the maximally mixed state and its minimum at pure states.
Entropy minimization under linear, spectral, or symmetry-induced constraints arises naturally in quantum statistical mechanics and quantum information theory. Classical examples include Gibbs states, microcanonical ensembles, and variational principles subject to conserved quantities or superselection rules. In such settings, entropy minimizers often admit an explicit description.
In practice, however, one rarely encounters exact minimizers. Instead, one typically works with states whose entropy is close, but not equal, to the minimum allowed by the constraints. This leads to the following natural stability question:
Does near-minimal entropy force quantitative closeness to an entropy minimizer?
This is a stability problem rather than a minimization problem: it asks for explicit control of distance to the minimizing set in terms of the entropy gap, rather than identification of the minimizers themselves. From an optimization viewpoint, the stability problem considered here is closely related to quadratic growth and error-bound properties for convex functionals. The result shows that, under a confining (fixed-support) constraint, the entropy functional exhibits quadratic growth relative to the set of minimizers when measured in trace norm. In this sense, the stability estimate can be interpreted as a second-order error bound for entropy minimization on a structured convex set, with constants determined explicitly by the geometry of the constraint.
1.2 Block-diagonal constraints
We focus on a particularly natural and flexible class of constraints. Suppose that
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathcal{H}_{i}$
is a fixed orthogonal decomposition into nontrivial subspaces. This decomposition induces a convex subset of $\mathcal{S}(\mathcal{H})$ consisting of states that are block-diagonal with respect to the given splitting.
Such block-diagonal constraints arise in a variety of standard settings, including:
  • symmetry reductions and superselection rules,
  • joint spectral decompositions of commuting observables,
  • coarse-graining procedures in quantum statistical mechanics,
  • measurement-induced decoherence.
A canonical example is provided by a unitary representation of a finite group $G$ on $\mathcal{H}$. States commuting with the representation decompose according to the isotypic components of $\mathcal{H}$, yielding a block-diagonal structure. Entropy minimization within this symmetry-restricted class selects states that are extremal relative to the representation, and stability quantifies how close a near-minimizer must be to such symmetry-pure states.
For block-diagonal states, the entropy admits a canonical decomposition into a classical (Shannon) contribution arising from block weights and a quantum contribution arising from the internal states within each block. This structural decomposition enables a precise analysis of entropy minimization and stability. Although the presentation uses quantum states for concreteness, no genuinely quantum phenomenon is invoked; the analysis applies verbatim to classical probability simplices and to finite-dimensional noncommutative state spaces.
The block structure considered here is not a technical artefact, but a concrete realisation of stratified convex geometry, which is the true object of the analysis.
1.3 Main result (informal)
Throughout this work, quadratic entropy–distance stability is established exclusively within the confining (fixed-support) regime. Boundary-transversal regimes, in which admissible perturbations explore directions transversal to the support of an entropy minimizer, exhibit a fundamentally different, non-quadratic behavior and are treated separately.
For a canonical class of block-diagonal constraints (see Section 6), the entropy minimization problem admits explicit solutions: in particular, the minimal entropy equals $\log r$ and is achieved by uniform mixtures of pure states, one from each block.
Here, dimension-independent means independent of the ambient Hilbert space dimension $\dim(\mathcal{H})$. The associated constants may depend on structural parameters fixed by the constraint geometry (e.g., the maximal active block dimension or minimal marginal weights), which are part of the model and remain uniform throughout the admissible set.
Our main result shows that this characterisation is quantitatively stable:
If a block-diagonal state has entropy at most $\log r+\varepsilon$, then it is within $O(\sqrt{\varepsilon})$ in trace norm of an entropy-minimizing state.
The proof yields explicit constants depending only on the geometry of the underlying convex constraint set, and not on the ambient Hilbert space dimension. Moreover, we show by explicit examples that the $\sqrt{\varepsilon}$ rate is optimal, even in purely classical (commutative) settings.
1.4 Related Work and Contribution
Our stability result is related to several classical inequalities in information theory and convex optimization.
Pinsker’s inequality [Pinsker1964, Csiszar1967] provides a fundamental bound relating relative entropy to trace distance:
(1) $D(\rho\|\sigma)\geq\frac{1}{2}\|\rho-\sigma\|_{1}^{2}.$
On fixed-support strata where $\mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma)$, the entropy gap is locally equivalent to the relative entropy $D(\rho\|\sigma)$ up to second order. In particular, Pinsker–type inequalities imply quadratic control of the trace distance in a neighbourhood of $\sigma$. In this sense, Theorem 4.1 is consistent with Pinsker’s inequality when a fixed reference state is present. The relationship between entropy differences and relative entropy has been further studied by Reeb [Reeb2015].
Classical continuity and metric inequalities for entropy include the works of Pinsker [Pinsker1964, Csiszar1967], Fannes [Fannes1973], Audenaert [Audenaert2007], Winter [Winter2016], and related developments in quantum information theory.
Entropy continuity has been extensively analyzed in the quantum information literature. Fannes [Fannes1973] established the first continuity bound for von Neumann entropy. Audenaert [Audenaert2007] provided sharp continuity estimates, and Winter [Winter2016] proved tight uniform continuity bounds valid across the entire state space. These results give global moduli of continuity for entropy as a function of states.
Our contribution relative to this prior work is threefold:
  1. 1.
    Geometric characterization of the quadratic regime: We provide a precise definition of the confining (fixed-support) regime (Definition 3.3) under which quadratic stability holds. This regime is characterized geometrically in terms of tangency to support strata and nondegeneracy of constrained entropy minimizers.
  2. 2.
    Explicit stability constants from constraint geometry: For block-separable constraint sets, we derive explicit formulas for the stability constant $C$ in Theorem 4.1. The constant decomposes into:
    • A marginal component $c_{1}$ determined by the Hessian of Shannon entropy $H(p)=-\sum_{i}p_{i}\log p_{i}$ restricted to faces of the marginal polytope $\Pi$,
    • A conditional component $c_{2}$ determined by the Hessian of von Neumann entropy $S(\rho_{i})$ at conditional entropy minimizers within each block $\mathcal{C}_{i}$.
    This geometric dependence on constraint curvature is made fully explicit in our analysis.
  3. 3.
    Optimality of the quadratic exponent: In Section 5, we prove that the exponent $2$ in the stability bound $S(\rho)-S_{\min}\geq C\|\rho-\sigma\|_{1}^{2}$ cannot be improved. We construct explicit families of states (both classical and quantum) for which the entropy gap scales precisely as $\varepsilon^{2}$ where $\varepsilon=\|\rho-\sigma\|_{1}$, precluding any uniform bound with exponent $\alpha<2$.
The block-separable structure studied here arises naturally in quantum systems with superselection rules, symmetry-reduced problems, and measurement-induced decoherence. Our analysis shows that this structure enables a complete decomposition of the stability problem into independent marginal and conditional subproblems.
Unlike Pinsker–type inequalities or entropy continuity bounds, our stability estimate does not compare two fixed states, nor does it rely on a reference distribution specified a priori. Instead, it controls the distance to an entropy minimizer that is defined intrinsically by the geometry of the constraint set itself and emerges only after solving the constrained variational problem.
1.5 Relation to existing work
Stability phenomena for entropy and related functionals have been studied in several contexts, including Pinsker–type inequalities for relative entropy, logarithmic Sobolev inequalities, and stability results in convex optimisation. These results typically control relative entropy with respect to a fixed reference state or quantify entropy decay along dynamical flows.
The present work differs in two essential respects:
  1. 1.
    We study entropy near its minimum as a variational problem on a constraint set, rather than entropy relative to a fixed reference state.
  2. 2.
    The constraint set is a nontrivial convex subset determined by block or spectral structure, rather than the full state space.
In particular, classical results such as Pinsker’s inequality do not imply the stability estimate proved here, since no reference state is fixed a priori and the minimizer emerges intrinsically from the geometry of the constraint set.
Optimality
The quadratic entropy–distance exponent established in this work is optimal. Even in purely classical (commutative) settings, no uniform linear or subquadratic stability bound can hold in a neighbourhood of entropy minimizers. This limitation is intrinsic and is demonstrated by explicit counterexamples in Section 5.
1.6 Operational significance and scope
Beyond the abstract variational analysis, it is important to indicate how the geometric hypotheses underlying the stability theorem manifest in concrete settings. Section 8 therefore presents a minimal finite-dimensional model, motivated by leakage phenomena in quantum information processing, which illustrates the practical relevance of the confining versus boundary–transversal distinction introduced above.
The purpose of this example is not to derive additional theoretical results, but to show that entropy-based certification is effective precisely in the confining (fixed-support) regime and necessarily fails once transversal leakage directions are permitted.
1.7 Structure of the paper
Section 2 introduces notation and recalls basic entropy inequalities. Section 3 analyses the geometry of block-diagonal convex constraint sets and characterises entropy minimizers. Section 4 establishes the main quantitative stability theorem with a complete proof. Section 5 proves optimality of the stability exponent through explicit classical and quantum examples. Section 6 discusses interpretations and consequences of the stability theorem. Section 7 relates the results to classical entropy inequalities. Section 8 provides computational guidance, showing how the marginal and conditional stability constants can be computed from Hessian eigenvalues for specific constraint geometries. The appendices collect standard entropy facts and technical arguments used throughout the paper.
2 Preliminaries and Notation
Throughout, $\mathcal{H}$ denotes a finite-dimensional complex Hilbert space of dimension $d\geq 1$. All operators are linear and act on $\mathcal{H}$ unless otherwise stated. Finite-dimensionality is assumed solely to ensure compactness and equivalence of norms; no dimension-dependent constants appear in the main results.
Finite dimensionality is assumed throughout and is essential for compactness and norm equivalence.
2.1 Quantum states and norms
Let $\mathcal{B}(\mathcal{H})$ denote the algebra of all linear operators on $\mathcal{H}$. A quantum state is an operator $\rho\in\mathcal{B}(\mathcal{H})$ satisfying
$\rho\geq 0,\qquad\operatorname{Tr}(\rho)=1.$
The set of all quantum states on $\mathcal{H}$ is denoted by $\mathcal{S}(\mathcal{H})$.
For $1\leq p<\infty$, we equip $\mathcal{B}(\mathcal{H})$ with the Schatten $p$-norms
$\|X\|_{p}:=\big(\operatorname{Tr}(|X|^{p})\big)^{1/p},\qquad|X|:=(X^{*}X)^{1/2}.$
Of particular importance is the trace norm $\|\cdot\|_{1}$, which metrizes statistical distinguishability of quantum states. For $\rho,\sigma\in\mathcal{S}(\mathcal{H})$, the trace distance $\|\rho-\sigma\|_{1}$ equals twice the maximal bias achievable in distinguishing $\rho$ from $\sigma$ via quantum measurements.
2.2 Von Neumann entropy and relative entropy
The von Neumann entropy [vonNeumann1927, NielsenChuang] of a state $\rho\in\mathcal{S}(\mathcal{H})$ is defined by
$S(\rho):=-\operatorname{Tr}(\rho\log\rho),$
where the logarithm is taken in the functional calculus sense, with the convention $0\log 0=0$. It satisfies the bounds
$0\leq S(\rho)\leq\log d,$
with equality on the left if and only if $\rho$ is pure, and equality on the right if and only if $\rho=\frac{1}{d}I$.
For $\rho,\sigma\in\mathcal{S}(\mathcal{H})$ with $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$, the quantum relative entropy is defined by
$D(\rho\|\sigma):=\operatorname{Tr}\!\big(\rho(\log\rho-\log\sigma)\big),$
and is set to $\infty$ otherwise. Relative entropy is nonnegative, jointly convex, and vanishes if and only if $\rho=\sigma$. It will be used only as a technical tool in the stability analysis.
2.3 Entropy inequalities
We recall a basic structural identity for block-diagonal states.
Lemma 2.1 (Entropy decomposition for block-diagonal states).
Let
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathcal{H}_{i}$
be an orthogonal decomposition, and let $\rho\in\mathcal{S}(\mathcal{H})$ be block-diagonal with respect to this decomposition:
$\rho=\bigoplus_{i=1}^{r}p_{i}\rho_{i},$
where $p_{i}\geq 0$, $\sum_{i}p_{i}=1$, and $\rho_{i}\in\mathcal{S}(\mathcal{H}_{i})$. Then
$S(\rho)=H(p)+\sum_{i=1}^{r}p_{i}S(\rho_{i}),$
where $H(p):=-\sum_{i}p_{i}\log p_{i}$ denotes the Shannon entropy.
Proof 2.2.
Since the blocks act on orthogonal subspaces, the spectrum of $\rho$ is the union of the spectra of the operators $p_{i}\rho_{i}$. The identity follows directly.
Lemma 2.3 (Pinsker inequality).
For all $\rho,\sigma\in\mathcal{S}(\mathcal{H})$,
$D(\rho\|\sigma)\geq\tfrac{1}{2}\|\rho-\sigma\|_{1}^{2}.$
2.4 Purity and effective dimension
The purity of a state $\rho$ is defined by
$\operatorname{Pur}(\rho):=\operatorname{Tr}(\rho^{2}).$
The purity satisfies $1/d\leq\operatorname{Pur}(\rho)\leq 1$.
We define the effective dimension by
$d_{\mathrm{eff}}(\rho):=\frac{1}{\operatorname{Pur}(\rho)}.$
Lemma 2.4 (Entropy–purity bound).
For all $\rho\in\mathcal{S}(\mathcal{H})$,
$S(\rho)\geq-\log\operatorname{Tr}(\rho^{2}).$
Proof 2.5.
This follows from Jensen’s inequality applied to the convex function $x\mapsto-\log x$ and the spectral decomposition of $\rho$.
2.5 Constrained entropy minimization
Let $\mathcal{C}\subset\mathcal{S}(\mathcal{H})$ be a closed convex subset. We define the minimal entropy on $\mathcal{C}$ by
$S_{\min}(\mathcal{C}):=\inf_{\rho\in\mathcal{C}}S(\rho).$
A state $\rho\in\mathcal{C}$ is called an entropy minimizer if $S(\rho)=S_{\min}(\mathcal{C})$. The central problem addressed here is the stability of entropy minimizers: given $\rho\in\mathcal{C}$ with $S(\rho)-S_{\min}(\mathcal{C})$ small, how close must $\rho$ be (in trace norm) to the set of minimizers?
Remark 2.6.
Entropy minimization is typically non-smooth and leads to highly structured extremal states. Quantitative stability therefore requires genuinely second-order control and cannot be obtained from first-order variational arguments alone.
3 Convex Block Constraints and Geometric Structure
This section introduces the class of constraint sets for which entropy minimization exhibits rigidity and quantitative stability. The framework is purely finite-dimensional and operator-theoretic.
3.1 Orthogonal block decompositions
Let $\mathcal{H}$ be a finite-dimensional Hilbert space and fix an orthogonal decomposition
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathcal{H}_{i}$
into nonzero subspaces. Let $P_{i}$ denote the orthogonal projection onto $\mathcal{H}_{i}$.
Every operator $X\in\mathcal{B}(\mathcal{H})$ admits the decomposition
$X=\sum_{i,j=1}^{r}P_{i}XP_{j}.$
Definition 3.1 (Block-diagonal algebra).
The block-diagonal algebra associated with the decomposition is
$\mathcal{A}:=\bigoplus_{i=1}^{r}\mathcal{B}(\mathcal{H}_{i}),$
consisting of all operators $X$ satisfying $P_{i}XP_{j}=0$ for $i\neq j$.
3.2 Block-diagonal states
Definition 3.2 (Block-diagonal states).
A state $\rho\in\mathcal{S}(\mathcal{H})$ is called block-diagonal (or block-commuting) if $\rho\in\mathcal{A}$, i.e.
$\rho=\sum_{i=1}^{r}P_{i}\rho P_{i}.$
Every block-diagonal state admits a unique representation
$\rho=\bigoplus_{i=1}^{r}p_{i}\rho_{i},$
where
$p_{i}:=\operatorname{Tr}(P_{i}\rho),\qquad\rho_{i}:=\frac{P_{i}\rho P_{i}}{p_{i}}\quad(p_{i}>0),$
and $\rho_{i}\in\mathcal{S}(\mathcal{H}_{i})$. If $p_{i}=0$, $\rho_{i}$ is arbitrary and plays no role.
3.3 Block-convex constraint sets
We now define the class of constraint sets considered here.
Definition 3.3 (Block-convex constraint set).
A subset $\mathcal{C}\subset\mathcal{S}(\mathcal{H})$ is called block-convex (with respect to the given decomposition) if there exist:
  • a compact convex set $\Pi\subset\Delta_{r}$,
  • compact convex sets $\mathcal{C}_{i}\subset\mathcal{S}(\mathcal{H}_{i})$,
such that
$\mathcal{C}=\left\{\bigoplus_{i=1}^{r}p_{i}\rho_{i}:p\in\Pi,\;\rho_{i}\in\mathcal{C}_{i}\right\}.$
Lemma 3.4 (Convexity of block-convex sets).
Every block-convex set $\mathcal{C}$ is compact and convex.
Proof 3.5.
Let $\rho=\oplus_{i}p_{i}\rho_{i}$ and $\sigma=\oplus_{i}q_{i}\sigma_{i}$ belong to $\mathcal{C}$, and let $0\leq\lambda\leq 1$. Then
$\lambda\rho+(1-\lambda)\sigma=\bigoplus_{i=1}^{r}\bigl(\lambda p_{i}+(1-\lambda)q_{i}\bigr)\,\tilde{\rho}_{i},$
where
$\tilde{\rho}_{i}=\frac{\lambda p_{i}\rho_{i}+(1-\lambda)q_{i}\sigma_{i}}{\lambda p_{i}+(1-\lambda)q_{i}}\quad\text{whenever }\lambda p_{i}+(1-\lambda)q_{i}>0.$
Since $\Pi$ and each $\mathcal{C}_{i}$ are convex, we have $\lambda p+(1-\lambda)q\in\Pi$ and $\tilde{\rho}_{i}\in\mathcal{C}_{i}$. Compactness follows from finite-dimensionality.
3.4 Entropy minimizers: marginal structure
Let $\mathcal{C}$ be block-convex and define
$S_{\min}:=\inf_{\rho\in\mathcal{C}}S(\rho).$
Lemma 3.6 (Extreme marginal property).
If $\rho=\oplus_{i}p_{i}\rho_{i}$ minimizes entropy over $\mathcal{C}$, then $p\in\mathrm{ext}(\Pi)$.
Proof 3.7.
By Lemma 2.1,
$S(\rho)=H(p)+\sum_{i=1}^{r}p_{i}S(\rho_{i}).$
The Shannon entropy $H$ is strictly concave on every face of $\Delta_{r}$. The second term is affine in $p$. If $p$ were not extreme in $\Pi$, then $p$ would be a nontrivial convex combination of distinct points in $\Pi$, strictly lowering $H(p)$ at one endpoint and contradicting minimality.
3.5 Entropy minimizers: conditional structure
Lemma 3.8 (Conditional minimization).
Let $\rho=\oplus_{i}p_{i}\rho_{i}$ be an entropy minimizer in $\mathcal{C}$. Then for every $i$ with $p_{i}>0$, the conditional state $\rho_{i}$ minimizes entropy over $\mathcal{C}_{i}$.
Proof 3.9.
Fix $i$ with $p_{i}>0$. Suppose $\rho_{i}$ is not an entropy minimizer in $\mathcal{C}_{i}$. Then there exists $\sigma_{i}\in\mathcal{C}_{i}$ with $S(\sigma_{i})<S(\rho_{i})$. Replacing $\rho_{i}$ by $\sigma_{i}$ while keeping all other blocks unchanged yields a state in $\mathcal{C}$ with strictly smaller entropy, contradicting minimality.
3.6 Structure of entropy minimizers
Proposition 3.10 (Structure of entropy minimizers).
Every entropy minimizer $\rho\in\mathcal{C}$ has the form
$\rho=\bigoplus_{i\in I}p_{i}\rho_{i},$
where:
  1. 1.
    $I\subset\{1,\dots,r\}$ is the minimal support of an extreme point $p\in\mathrm{ext}(\Pi)$,
  2. 2.
    $p_{i}>0$ for $i\in I$ and $p_{i}=0$ otherwise,
  3. 3.
    each $\rho_{i}$ is an entropy minimizer in $\mathcal{C}_{i}$.
Remark 3.11.
The index set $I$ is uniquely determined by the support of the extreme marginal $p$. No minimality assumption beyond this support condition is required.
3.7 Compactness of the minimizer set
Let $\mathcal{M}\subset\mathcal{C}$ denote the set of entropy minimizers.
Lemma 3.12 (Compactness).
The set $\mathcal{M}$ is compact.
Proof 3.13.
The set $\mathcal{C}$ is compact, and entropy is lower semicontinuous. Hence the set of minimizers is closed and therefore compact.
3.8 Preliminary continuity estimate
The following qualitative statement will be sharpened in Section 4.
Lemma 3.14 (Qualitative stability).
For every $\varepsilon>0$ there exists $\delta>0$ such that
$S(\rho)\leq S_{\min}+\delta\quad\Longrightarrow\quad\operatorname{dist}_{1}(\rho,\mathcal{M})\leq\varepsilon.$
Proof 3.15.
This follows from compactness of $\mathcal{M}$ and lower semicontinuity of entropy.
4 A Quantitative Stability Theorem
This section establishes the main result: entropy minimizers under block-convex constraints are quantitatively rigid. We show that entropy excess controls the squared trace-norm distance to the minimizer set, with constants depending only on the geometry of the constraint. Before proceeding to the local second-order analysis, we note that by continuity of the entropy functional on the compact constraint set and compactness of the set of entropy minimizers, there exists $\varepsilon_{0}>0$ such that any $\rho\in\mathcal{C}$ satisfying $S(\rho)-S_{\min}\leq\varepsilon_{0}$ lies in a neighbourhood of some entropy minimizer $\sigma\in\mathcal{M}$ where the fixed-support hypothesis holds and the local Taylor expansion applies.
4.1 Setup and notation
Let $\mathcal{C}\subset\mathcal{S}(\mathcal{H})$ be a block-convex constraint set as in Definition 3.3. Define the minimal entropy
$S_{\min}:=\inf_{\rho\in\mathcal{C}}S(\rho),$
and let $\mathcal{M}\subset\mathcal{C}$ denote the (compact) set of entropy minimizers.
Every $\rho\in\mathcal{C}$ admits a decomposition
$\rho=\bigoplus_{i=1}^{r}p_{i}\rho_{i},$
with $p\in\Pi$ and $\rho_{i}\in\mathcal{C}_{i}$. For $\sigma\in\mathcal{M}$, we write
$\sigma=\bigoplus_{i=1}^{r}q_{i}\sigma_{i}.$
Throughout, the marginal polytope $\Pi$ is assumed to be a compact polytope with finitely many faces.
4.2 Statement of the stability theorem
Throughout, the confining (fixed-support) hypothesis is understood uniformly in a neighborhood of the minimizer set $\mathcal{M}$.
Theorem 4.1 (Entropy stability under block-convex constraints).
Assume that the marginal minimizer satisfies Assumption 1 and that all entropy minimizers $\sigma\in\mathcal{M}$ satisfy the confining (fixed-support) hypothesis of Lemma 4.7.
Let $\mathcal{C}\subset\mathcal{S}(\mathcal{H})$ be a block–convex constraint set in the sense of Definition 3.3, induced by a fixed block decomposition of $\mathcal{H}$, and let $\mathcal{M}$ denote the set of entropy minimizers on $\mathcal{C}$.
Then there exists a constant $C=C(\Pi,\{\mathcal{C}_{i}\})>0$ such that for all $\rho\in\mathcal{C}$,
$S(\rho)-S_{\min}\;\geq\;C\,\operatorname{dist}_{1}(\rho,\mathcal{M})^{2},$
where
$\operatorname{dist}_{1}(\rho,\mathcal{M}):=\inf_{\sigma\in\mathcal{M}}\|\rho-\sigma\|_{1}.$
Remark 4.2 (Scope and interpretation).
The theorem establishes quadratic entropy stability in the fixed-support (confining) regime. When constrained entropy minimizers lie on the boundary of state space and admissible perturbations explore directions transversal to the support, the quadratic stability mechanism used here necessarily fails or requires modification. Such boundary–transversal regimes are not addressed in the present work.
From an optimization viewpoint, Theorem 4.1 establishes a quadratic growth (error bound) property for the entropy functional restricted to a structured convex constraint set. The stability constant $C=C(\Pi,\{\mathcal{C}_{i}\})$ depends only on the geometry of the marginal polytope and the conditional constraint sets; an explicit form is discussed in Section 4.8.
Dependence of constants
The constant $C$ may depend on $d_{\max}:=\max_{i}\dim(H_{i})$ (or, equivalently, on the maximal active support rank imposed by the constraint geometry), but is independent of the ambient Hilbert space dimension once the block structure and confinement parameters are fixed.
4.3 Reduction to a fixed minimizer
Fix $\rho\in\mathcal{C}$ and choose $\sigma\in\mathcal{M}$ such that
$\|\rho-\sigma\|_{1}=\operatorname{dist}_{1}(\rho,\mathcal{M}).$
By Proposition 3.10, the marginal $q$ is an extreme point of $\Pi$, and for each $i$ with $q_{i}>0$, $\sigma_{i}$ minimizes entropy on $\mathcal{C}_{i}$.
4.4 Marginal entropy stability
We first control entropy increase arising from deviation of the marginal distribution $p$ from the extreme point $q$.
Assumption 1 (Marginal confinement / fixed support in $\Pi$).
Let $q\in\operatorname{ext}(\Pi)$ be an entropy-minimizing marginal.
There exist an index set $I\subset\{1,\dots,r\}$ and a constant $\underline{q}>0$ such that:
  1. 1.
    $\operatorname{supp}(q)=I$ and $q_{i}\geq\underline{q}$ for all $i\in I$;
  2. 2.
    all admissible marginals $p\in\Pi$ in a neighborhood of $q$ satisfy $\operatorname{supp}(p)=I$ (equivalently, $\Pi$ is locally contained in the face $\Delta_{I}$).
Lemma 4.3 (Marginal entropy rigidity).
Assume Assumption 1. Let $\Pi\subset\Delta_{r}$ be a compact convex polytope and let $q\in\mathrm{ext}(\Pi)$. Then there exists a constant $c_{1}>0$, depending only on $\Pi$, such that for all $p\in\Pi$,
$H(p)-H(q)\;\geq\;c_{1}\,\|p-q\|_{1}^{2}.$
Proof 4.4.
Let $F$ be the minimal face of $\Pi$ containing $q$. Restricted to the relative interior of $F$, the Shannon entropy H is twice continuously differentiable and strictly concave. Consequently, the Hessian of H-H restricted to the tangent space of $F$ at $q$ is positive definite. Therefore, there exists a constant $\lambda_{F}>0$ such that
$H(p)-H(q)\geq\lambda_{F}\,\|p-q\|_{2}^{2}\quad\text{for all }p\in F.$
Since $\Pi$ has finitely many faces, the constants $\lambda_{F}$ obtained in this way admit a strictly positive minimum over all faces of $\Pi$. Moreover, all norms are equivalent on the finite-dimensional affine span of $\Pi$. Combining these observations yields a uniform constant $c_{1}>0$ such that
$H(p)-H(q)\geq c_{1}\,\|p-q\|_{1}^{2}\quad\text{for all }p\in\Pi.$
Remark 4.5 (Local-to-global reduction for the marginal term).
The estimate in Lemma 4.3 is only used in the neighbourhood of the entropy-minimising extreme marginal $q$ selected in Section 4.3. If $\|p-q\|_{1}$ is bounded from below by a fixed constant, then H(p)H(q)H(p)-H(q) admits a strictly positive lower bound by compactness of $\Pi$, so the final stability inequality remains global after possibly decreasing the constant. Thus no loss of generality is incurred by interpreting Lemma 4.3 as a local rigidity statement around $q$.
4.5 Conditional entropy stability
Next we control entropy increase arising within each block.
Definition 4.6 (Nondegenerate constrained minimizer).
Let $C\subset S(H)$ be a constraint set and $\sigma\in C$. We say that $\sigma$ is a nondegenerate constrained minimizer of f:=Sf:=-S on the fixed-support stratum if there exists $m>0$ such that, for all admissible tangent directions $X$ to $C\cap D(PH)$ at $\sigma$ (with $P=\operatorname{supp}(\sigma)$), one has
$\langle X,\nabla^{2}f(\sigma)[X]\rangle\geq m\|X\|_{2}^{2}.$
where $\|\cdot\|_{2}$ denotes the Hilbert–Schmidt norm.
Lemma 4.7 (Entropy stability under confining constraints).
Let $\mathcal{C}$ be a convex constraint set of quantum states on a finite-dimensional Hilbert space $\mathcal{H}$, and let $\sigma\in\mathcal{C}$ be an entropy minimizer.
Assume that there exists a neighbourhood $U$ of $\sigma$ such that all admissible perturbations preserve the support of $\sigma$, i.e.
$P(\rho-\sigma)Q=Q(\rho-\sigma)P=0\qquad\text{for all }\rho\in\mathcal{C}\cap U,$
where $P$ is the support projection of $\sigma$ and Q=IPQ=I-P.
Assume moreover that $\sigma$ is a nondegenerate constrained minimizer of S-S on this fixed-support stratum.
Then there exist constants $c>0$ and $\varepsilon_{0}>0$ such that for all $\rho\in\mathcal{C}$ with
$S(\rho)-S(\sigma)\leq\varepsilon_{0},$
one has
$S(\rho)-S(\sigma)\ \geq\ c\,\|\rho-\sigma\|_{1}^{2}.$
Equivalently,
$\|\rho-\sigma\|_{1}\ \leq\ C\,\sqrt{S(\rho)-S(\sigma)},$
for some constant $C>0$.
Remark 4.8 (Scope of Lemma 4.7).
Lemma 4.7 applies to the fixed-support (confining) regime, that is, to situations in which admissible perturbations of the constrained entropy minimizer remain within a fixed-rank support stratum.
Proof 4.9.
By the confining (fixed-support) hypothesis, there exists a neighbourhood $U$ of $\sigma$ such that every admissible $\rho\in\mathcal{C}\cap U$ satisfies
$Q\rho=\rho Q=0,\qquad P=\operatorname{supp}(\sigma),\quad Q=I-P.$
Hence all admissible states lie in the fixed-support stratum
$\mathcal{S}(P\mathcal{H}):=\{\rho\in\mathcal{S}(\mathcal{H}):Q\rho=\rho Q=0\},$
on which $\sigma$ is faithful, i.e. $P\sigma P\succ 0$ as an operator on $P\mathcal{H}$.
On this stratum the von Neumann entropy is $C^{2}$ in a neighbourhood of $\sigma_{i}$, and f:=Sf:=-S has a nondegenerate constrained minimum at $\sigma_{i}$ by assumption. Therefore, in a local chart of the admissible set $C_{i}\cap U$ around $\sigma_{i}$, the restricted Hessian satisfies
$\nabla^{2}f(\sigma_{i})\ \succeq\ m_{i}\,I$
on the tangent space of $C_{i}$ at $\sigma_{i}$, for some $m_{i}>0$. A second-order Taylor expansion then yields, for $\rho_{i}$ sufficiently close to $\sigma_{i}$,
$f(\rho_{i})-f(\sigma_{i})\ \geq\ \frac{m_{i}}{2}\,\|\rho_{i}-\sigma_{i}\|_{2}^{2},$
and hence
$S(\rho_{i})-S(\sigma_{i})=f(\sigma_{i})-f(\rho_{i})\ \geq\ \frac{m_{i}}{2}\,\|\rho_{i}-\sigma_{i}\|_{2}^{2}.$
Finally, using $\|X\|_{2}\geq\|X\|_{1}/\sqrt{\dim(P_{i}\mathcal{H}_{i})}$, we obtain
$S(\rho_{i})-S(\sigma_{i})\ \geq\ \frac{m_{i}}{2\,\dim(P_{i}\mathcal{H}_{i})}\,\|\rho_{i}-\sigma_{i}\|_{1}^{2},$
which proves the claimed quadratic stability bound. Since the support projection $P$ and the admissible block dimensions are fixed by the constraint geometry, the resulting constants depend only on the constraint set and not on the ambient Hilbert space dimension.
Remark 4.10 (Why this excludes the boundary–transversal regime).
The argument above crucially uses the fixed-support (confining) hypothesis. In the boundary–transversal regime, admissible perturbations necessarily generate support–kernel coupling, and the entropy variation acquires an unavoidable $\delta^{2}\log(1/\delta)$ correction.
4.6 Assembly of the entropy estimate
Using entropy decomposition, we write
$S(\rho)-S(\sigma)=H(p)-H(q)+\sum_{i=1}^{r}p_{i}\bigl(S(\rho_{i})-S(\sigma_{i})\bigr).$
We note that any affine contribution of the form $\sum_{i}(p_{i}-q_{i})S(\sigma_{i})$ depends only on the fixed entropy-minimizing blocks $\sigma_{i}$ and may therefore be absorbed into the marginal entropy difference H(p)H(q)H(p)-H(q). This term does not affect the stability estimate.
Applying Lemmas 4.3 and 4.7 to the marginal and conditional contributions, respectively, we obtain the quantitative lower bound
$S(\rho)-S_{\min}\;\geq\;c_{1}\|p-q\|_{1}^{2}+\tfrac{1}{2}\sum_{i}p_{i}\|\rho_{i}-\sigma_{i}\|_{1}^{2}.$
4.7 Blockwise trace norm control
Lemma 4.11 (Blockwise trace inequality).
For block-diagonal states $\rho=\oplus_{i}p_{i}\rho_{i}$ and $\sigma=\oplus_{i}q_{i}\sigma_{i}$,
$\|\rho-\sigma\|_{1}^{2}\;\leq\;2\|p-q\|_{1}^{2}+2\sum_{i}p_{i}\|\rho_{i}-\sigma_{i}\|_{1}^{2}.$
Proof 4.12.
We decompose
$\rho-\sigma=\bigoplus_{i}(p_{i}-q_{i})\sigma_{i}+\bigoplus_{i}p_{i}(\rho_{i}-\sigma_{i}).$
Orthogonality of blocks and the triangle inequality give
$\|\rho-\sigma\|_{1}\leq\|p-q\|_{1}+\sum_{i}p_{i}\|\rho_{i}-\sigma_{i}\|_{1}.$
Squaring and applying $(a+b)^{2}\leq 2a^{2}+2b^{2}$ yields the claim.
4.8 Conclusion of the proof
This section makes explicit the dependence of the stability constant $C=C(\Pi,\{\mathcal{C}_{i}\})$ appearing in Theorem 4.1, by expressing it in terms of the marginal and conditional curvature constants introduced in the preceding lemmas.
Combining the previous inequalities, we obtain
$S(\rho)-S_{\min}\;\geq\;C\,\|\rho-\sigma\|_{1}^{2},$
with
$C=\frac{1}{2}\min\{c_{1},\tfrac{1}{2}\}.$
Since $M$ is compact and the confining/nondegeneracy hypotheses hold uniformly along $M$ by assumption, the local constants produced by Lemmas 4.34.7 admit a strictly positive global minimum over a finite subcover of $M$.
This completes the proof of Theorem 4.1.
5 Sharpness and Optimality of the Stability Exponent
In this section we show that the quadratic exponent in Theorem 4.1 is optimal. More precisely, we construct explicit families of states within block-convex constraint sets for which the entropy gap scales quadratically with the trace-norm distance to the set of minimizers. As a consequence, no stability inequality with exponent strictly less than 2 can hold uniformly under the stated hypotheses.
All examples are finite-dimensional and require no additional structure beyond that developed in Sections 3-4.
5.1 The classical simplex model
We begin with the purely classical case, which already captures the essential geometry.
Let $\mathcal{H}=\mathbb{C}^{r}$, and identify $\mathcal{S}(\mathcal{H})$ with the probability simplex $\Delta_{r}$. Let $\Pi\subset\Delta_{r}$ be a compact convex polytope, and define the constraint set
$\mathcal{C}:=\Pi.$
Entropy reduces to Shannon entropy $S(\rho)=H(p)$.
Let $q\in\mathrm{ext}(\Pi)$ be an extreme point. Then $q$ is an entropy minimizer on $\mathcal{C}$.
Proposition 5.1 (Quadratic sharpness in the classical case).
There exists a constant $c>0$ and a family $\{p_{\varepsilon}\}_{\varepsilon>0}\subset\Pi$ such that
$\|p_{\varepsilon}-q\|_{1}\sim\varepsilon,\qquad H(p_{\varepsilon})-H(q)\sim c\,\varepsilon^{2}\quad\text{as }\varepsilon\to 0.$
In particular, no bound of the form
$H(p)-H(q)\geq C\,\|p-q\|_{1}^{\alpha}$
can hold uniformly for any $\alpha<2$.
Proof 5.2.
Let $F$ be a face of $\Pi$ of dimension $k\geq 1$ containing $q$, and let $v\in\mathbb{R}^{r}$ be a tangent vector to $F$ at $q$ satisfying $\sum_{i}v_{i}=0$ and $q_{i}+tv_{i}\geq 0$ for small $t$.
Define
$p_{\varepsilon}:=q+\varepsilon v.$
Since $v$ is tangent to $F$, we have $p_{\varepsilon}\in\Pi$ for all sufficiently small $\varepsilon$.
Because $q$ is a minimizer of $H$ on $\Pi$, the first-order directional derivative of $H$ along $v$ vanishes:
$\left.\frac{d}{dt}H(q+tv)\right|_{t=0}=0.$
A second-order Taylor expansion yields
$H(p_{\varepsilon})-H(q)=\tfrac{1}{2}\,\varepsilon^{2}\,v^{\top}\bigl(\nabla^{2}H(q)\bigr)v+o(\varepsilon^{2}).$
Since $H$ is strictly concave on the relative interior of the face $F$, the Hessian of $H$ is negative definite on the tangent space of $F$ at $q$, implying the stated asymptotic behavior.
This shows that quadratic behavior is intrinsic to entropy near boundary minimizers, even in the absence of quantum structure.
5.2 Quantum block-diagonal models
We now lift the previous construction to genuinely quantum settings.
Let
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathbb{C}^{d_{i}},$
and let $\mathcal{C}$ consist of all block-diagonal states
$\rho=\bigoplus_{i=1}^{r}p_{i}\rho_{i},\qquad p\in\Pi,\;\rho_{i}\in\mathcal{S}(\mathbb{C}^{d_{i}}).$
Proposition 5.3 (Entropy minimizers).
Entropy minimizers in $\mathcal{C}$ are precisely the states
$\sigma=\bigoplus_{i=1}^{r}q_{i}\sigma_{i},$
where $q\in\mathrm{ext}(\Pi)$ and each $\sigma_{i}$ is pure.
Proof 5.4.
Entropy decomposes as
$S(\rho)=H(p)+\sum_{i}p_{i}S(\rho_{i}).$
Minimization forces $p=q$ extreme and each $S(\rho_{i})=0$, hence $\rho_{i}$ pure.
5.3 Sharpness under block perturbations
Fix a minimizer $\sigma=\bigoplus_{i}q_{i}\sigma_{i}$ as above, and let $p_{\varepsilon}$ be the marginal perturbation constructed in Proposition 5.1. Define
$\rho_{\varepsilon}:=\bigoplus_{i=1}^{r}p_{\varepsilon}(i)\,\sigma_{i}.$
Proposition 5.5 (Quadratic entropy growth).
The family $\{\rho_{\varepsilon}\}$ satisfies
$\|\rho_{\varepsilon}-\sigma\|_{1}\sim\varepsilon,\qquad S(\rho_{\varepsilon})-S(\sigma)\sim c\,\varepsilon^{2}\quad\text{as }\varepsilon\to 0.$
Proof 5.6.
Since the conditional states are unchanged, entropy variation is purely classical:
$S(\rho_{\varepsilon})-S(\sigma)=H(p_{\varepsilon})-H(q).$
The trace norm satisfies
$\|\rho_{\varepsilon}-\sigma\|_{1}=\|p_{\varepsilon}-q\|_{1},$
since blocks are orthogonal. The claim follows directly from Proposition 5.1.
5.4 Failure of linear stability
We summarize the sharpness result.
Theorem 5.7 (Optimality of the stability exponent).
Under the hypotheses of Theorem 4.1, the exponent $2$ is optimal. In general, there exists no constant $C>0$ and no exponent $\alpha<2$ such that
$S(\rho)-S_{\min}\geq C\,\operatorname{dist}_{1}(\rho,\mathcal{M})^{\alpha}$
holds uniformly for all $\rho\in\mathcal{C}$.
Proof 5.8.
The families constructed above satisfy
$S(\rho_{\varepsilon})-S_{\min}\sim c\,\|\rho_{\varepsilon}-\sigma\|_{1}^{2},$
which precludes any linear lower bound.
5.5 Interpretation
The obstruction to linear stability arises from flat directions along faces of the marginal polytope $\Pi$. Entropy exhibits genuine curvature only transverse to these faces, leading to quadratic - but not linear - rigidity.
This mechanism is intrinsic and persists in both classical and quantum settings.
6 Example: Pure State Entropy Minimization
In this section we provide a concrete example where the stability constants in Theorem 4.1 can be computed explicitly. The constraint set fixes block populations while allowing states within each block to vary freely, yielding entropy minimizers that are pure states in each block. This example illustrates the decomposition of the stability constant into marginal and conditional components and demonstrates the sharpness of the quadratic bound.
6.1 Setup
Let $\mathcal{H}$ be a finite-dimensional Hilbert space and suppose
$\mathcal{H}=\bigoplus_{i=1}^{r}\mathcal{H}_{i}$
is a fixed orthogonal decomposition into nontrivial subspaces, with $P_{i}$ denoting the orthogonal projection onto $\mathcal{H}_{i}$.
Define the block-diagonal state space
$\mathcal{S}_{\mathrm{bd}}(\mathcal{H})=\left\{\rho\in\mathcal{S}(\mathcal{H}):\rho=\sum_{i=1}^{r}P_{i}\rho P_{i}\right\}.$
Such decompositions arise naturally from symmetry reductions, superselection rules, or coarse-grained descriptions of larger systems.
6.2 Fixed Block Populations
Fix a probability vector $q=(q_{1},\ldots,q_{r})$ with $q_{i}>0$ and $\sum_{i}q_{i}=1$. Consider the affine constraint set
$\mathcal{C}_{q}:=\left\{\rho\in\mathcal{S}_{\mathrm{bd}}(\mathcal{H}):\mathrm{Tr}(P_{i}\rho)=q_{i}\text{ for all }i\right\}.$
This constraint fixes the population (total probability) in each block while allowing the internal state within each block to vary. The set $\mathcal{C}_{q}$ is block-convex in the sense of Definition 3.3, with $\Pi=\{q\}$ (a single point) and $\mathcal{C}_{i}=\mathcal{S}(\mathcal{H}_{i})$ (the full state space of block $i$).
Proposition 6.1 (Entropy minimizers).
The set of entropy minimizers in $\mathcal{C}_{q}$ consists precisely of the states
$\sigma_{q}=\bigoplus_{i=1}^{r}q_{i}\,|\psi_{i}\rangle\langle\psi_{i}|,$
where each $|\psi_{i}\rangle$ is an arbitrary unit vector in $\mathcal{H}_{i}$. Equivalently, the minimizers are product states with pure states in each block, unique up to unitary transformations within each block.
Proof 6.2.
For $\rho=\bigoplus_{i}q_{i}\rho_{i}\in\mathcal{C}_{q}$, entropy decomposes as
$S(\rho)=H(q)+\sum_{i}q_{i}S(\rho_{i}).$
Since $H(q)$ is constant on $\mathcal{C}_{q}$, entropy is minimized by minimizing each $S(\rho_{i})$. The von Neumann entropy $S(\rho_{i})$ is minimized when $\rho_{i}$ is a pure state (rank-one projection), giving $S(\rho_{i})=0$.
6.3 Quantitative Stability Estimate
We now derive an explicit quantitative stability estimate in this setting.
Theorem 6.3 (Stability for fixed block populations).
Let $q=(q_{1},\dots,q_{r})$ with $q_{i}>0$ and $\sum_{i}q_{i}=1$, and let $\mathcal{C}_{q}$ be as above. Let
$\mathcal{M}_{q}:=\left\{\bigoplus_{i=1}^{r}q_{i}\,|\psi_{i}\rangle\langle\psi_{i}|:\ |\psi_{i}\rangle\in\mathcal{H}_{i},\ \|\psi_{i}\|=1\right\}$
denote the set of entropy minimizers in $\mathcal{C}_{q}$, and write $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q}):=\inf_{\sigma\in\mathcal{M}_{q}}\|\rho-\sigma\|_{1}$.
Then for every $\rho=\bigoplus_{i}q_{i}\rho_{i}\in\mathcal{C}_{q}$, the entropy gap satisfies
$S(\rho)-S_{\min}=\sum_{i=1}^{r}q_{i}\,S(\rho_{i}),\qquad S_{\min}=H(q),$
and the following quantitative implication holds: if $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\leq 1$, then
$S(\rho)-S_{\min}\;\geq\;\frac{1}{2}\,\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})^{2}.$
Equivalently, whenever $S(\rho)-S_{\min}\leq\frac{1}{2}$, one has
$\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\;\leq\;\sqrt{2\,(S(\rho)-S_{\min})}.$
Proof 6.4.
The entropy decomposition follows from block-diagonality: for $\rho=\bigoplus_{i}q_{i}\rho_{i}$, one has $S(\rho)=H(q)+\sum_{i}q_{i}S(\rho_{i})$.
For each block $i$, define the distance to the set of pure states:
$d_{i}:=\mathrm{dist}_{1}\bigl(\rho_{i},\mathrm{Pure}(\mathcal{H}_{i})\bigr)=\inf_{\|\psi\|=1}\|\rho_{i}-|\psi\rangle\langle\psi|\|_{1}.$
By trace-norm additivity over orthogonal blocks,
(2) $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})=\sum_{i=1}^{r}q_{i}\,d_{i}.$
Let $\lambda_{i}:=\lambda_{\max}(\rho_{i})$ be the largest eigenvalue of $\rho_{i}$. Choosing $|\psi_{i}\rangle$ to be an eigenvector corresponding to $\lambda_{i}$ gives $d_{i}=2(1-\lambda_{i})$.
By Schur concavity of entropy, among density matrices with fixed largest eigenvalue $\lambda_{i}$, the entropy $S(\rho_{i})$ is minimized when the remaining probability mass $1-\lambda_{i}$ is concentrated on a single orthogonal eigenvector. This gives the lower bound
$S(\rho_{i})\ \geq\ h(1-\lambda_{i}),\qquad h(t):=-(1-t)\log(1-t)-t\log t,$
where $h$ is the binary entropy function.
If $d_{i}\leq 1$, then $t_{i}:=1-\lambda_{i}=d_{i}/2\in[0,\frac{1}{2}]$. On this interval, $h$ satisfies the quadratic lower bound
[…]
which follows from $h(0)=h^{\prime}(0)=0$ and $h^{\prime\prime}(t)=\frac{1}{t(1-t)}\geq 4$ on $0,\frac{1}{2}]$.
Therefore, whenever $d_{i}\leq 1$,
$S(\rho_{i})\ \geq\ \frac{1}{2}\,d_{i}^{2}.$
Assuming $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\leq 1$, we have $\sum_{i}q_{i}d_{i}\leq 1$, which implies $d_{i}\leq 1$ for all $i$. Summing the previous inequality with weights $q_{i}$ yields
$S(\rho)-S_{\min}=\sum_{i}q_{i}S(\rho_{i})\ \geq\ \frac{1}{2}\sum_{i}q_{i}d_{i}^{2}\ \geq\ \frac{1}{2}\left(\sum_{i}q_{i}d_{i}\right)^{2},$
where the last step uses Jensen’s inequality for the convex function $x\mapsto x^{2}$. Using (2) gives the claimed bound. The final implication is immediate.
6.4 Explicit Constant in the Uniform Case
The stability constant can be computed explicitly when all block populations are equal.
Assume that $q_{i}=\frac{1}{r}$ for all $i$. Then the entropy minimizers are
$\sigma_{q}=\frac{1}{r}\bigoplus_{i=1}^{r}|\psi_{i}\rangle\langle\psi_{i}|.$
Corollary 6.5 (Uniform block populations).
Assume that $q_{i}=\frac{1}{r}$ for all $i$. Then for all $\rho\in\mathcal{C}_{q}$ with $\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\leq 1$,
$S(\rho)-S_{\min}\;\geq\;\frac{1}{2}\,\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})^{2}.$
In particular, if $S(\rho)-S_{\min}\leq\frac{1}{2}$, then
$\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\;\leq\;\sqrt{2\,(S(\rho)-S_{\min})}.$
Proof 6.6.
This follows immediately from Theorem 6.3 specialized to the uniform marginal distribution $q$.
6.5 Optimality of the Quadratic Exponent
The inequality in Theorem 6.3 gives
$\mathrm{dist}_{1}(\rho,\mathcal{M}_{q})\;\leq\;\sqrt{2\,(S(\rho)-S_{\min})}.$
Thus, an entropy gap of order $\varepsilon$ implies the state is within $O(\sqrt{\varepsilon})$ in trace distance from the set of entropy minimizers. By Section 5, this quadratic exponent is optimal: no linear bound of the form
$\|\rho-\sigma_{q}\|_{1}\leq C\,(S(\rho)-S_{\min})$
can hold uniformly, even in this simple example.
This demonstrates the sharpness of the quadratic stability principle and confirms that the $\sqrt{\varepsilon}$ distance scaling cannot be improved without additional structure.
6.6 Relation to Standard Entropy Inequalities
This stability result is not a direct consequence of standard entropy inequalities:
  • Pinsker’s inequality $D(\rho\|\sigma)\geq\frac{1}{2}\|\rho-\sigma\|_{1}^{2}$ provides a lower bound relating relative entropy to trace distance. In our setting, the reference state $\sigma_{q}$ emerges from constrained optimization rather than being fixed a priori. The constraint structure (fixed block populations) enables explicit computation of the stability constant.
  • Entropy continuity bounds (Fannes, Audenaert, Winter) provide global moduli of continuity for von Neumann entropy. Our result is a local stability estimate near constrained minimizers, exploiting the specific geometry of block-separable constraints to obtain sharp constants.
  • Strong convexity: The von Neumann entropy is globally concave but not strongly concave. Quadratic stability emerges after restriction to the constraint set $\mathcal{C}_{q}$ and localization near minimizers in the fixed-support regime.
The explicit computation of the stability constant $C=\frac{1}{2}$ in Theorem 6.3 provides quantitative information beyond what standard entropy inequalities yield for this constrained optimization problem.
7 Relation to Classical Entropy Inequalities
The stability inequality established in Theorem 4.1 interacts with, but is fundamentally distinct from, several well-known entropy inequalities in information theory and convex optimization. We briefly clarify these relationships and emphasize the nonstandard features of the present result.
7.1 Relation to Pinsker, continuity, and functional inequalities
Pinsker’s inequality [Pinsker1964, Csiszar1967] provides a quantitative relation between relative entropy and trace distance by comparing two fixed states. Related continuity bounds for the von Neumann entropy were established by Fannes [Fannes1973], Audenaert [Audenaert2007], and Winter [Winter2016], yielding sharp global moduli of continuity for entropy as a function of states.
These inequalities are complementary in spirit to the present work: they control either relative entropy with respect to a fixed reference state or uniform continuity of entropy on the full state space. By contrast, Theorem 4.1 treats the entropy functional itself as a variational object over a structured constraint set, with no reference state fixed a priori; the entropy minimizer instead emerges intrinsically from the geometry of the constraint.
Log-Sobolev inequalities relate entropy to Dirichlet forms associated with Markov semigroups and play a central role in entropy decay and mixing theory [KastoryanoTemme2013, CarlenMaas2017, Brannan2022]. The present stability estimate is orthogonal in nature: it is purely static and geometric, involving neither dynamics nor a prescribed equilibrium state. Related perspectives on entropy inequalities and stability appear in classical convex analysis and information theory; see, e.g., [Kullback1959, Rockafellar1970, BonnansShapiro2000].
7.2 Quadratic growth and nonstandard features
It is well known that the von Neumann entropy is concave but not strongly concave on the full state space. Consequently, no global quadratic growth inequality can hold without further structural assumptions.
The nonstandard feature of Theorem 4.1 is that quadratic entropy–distance stability emerges locally near constrained minimizers after restriction to block-convex constraint sets. The entropy functional lacks global strong convexity, and the admissible set is stratified; quadratic growth therefore does not hold a priori, but arises only after confinement to a fixed-support stratum determined by the constraint geometry.
From an optimization viewpoint, the theorem establishes a quadratic growth (error bound) property for entropy minimization on a structured convex set, with constants depending only on the geometry of the constraint and not on the ambient Hilbert space dimension.
This setting differs fundamentally from classical majorization or continuity results, which describe qualitative extremality but do not provide quantitative stability estimates near entropy minimizers.
8 Computational remarks
The stability constant $C$ in Theorem 4.1 can be made explicit for concrete constraint sets by combining marginal and conditional second-order information.
Marginal contribution.
The marginal constant $c_{1}$ from Lemma 4.3 is determined by the strict concavity of Shannon entropy on faces of the marginal polytope $\Pi$. For an extreme point $q\in\operatorname{ext}(\Pi)$, let $F$ denote the minimal face containing $q$. The Hessian of $H(p)=-\sum_{i}p_{i}\log p_{i}$ restricted to the tangent space of $F$ at $q$ is negative definite, with eigenvalues given by $1/q_{i}$ along admissible directions. The constant $c_{1}$ is obtained by minimizing the smallest absolute eigenvalue over all such faces $F$. For polytopes specified by a finite block structure, this quantity can be computed explicitly or bounded from below.
Conditional contribution.
The conditional constant $c_{2}$ from Lemma 3.8 depends on the Hessian of von Neumann entropy at the conditional minimizers $\sigma_{i}\in\mathcal{C}_{i}$. Under the confining hypothesis, the entropy Hessian is uniformly positive on admissible tangent directions, with a lower bound $m_{i}>0$ determined by the local spectral gap and geometry of $\mathcal{C}_{i}$. Setting $c_{2}=\min_{i:q_{i}>0}m_{i}$, the overall stability constant takes the form
$C=\min\!\left\{c_{1},\;\frac{c_{2}}{2d_{\max}}\right\},$
where $d_{\max}$ accounts for norm equivalence on conditional subspaces.
Remarks on computation.
For constraint sets of moderate size, the constants $c_{1}$ and $c_{2}$ can be evaluated numerically using standard eigenvalue routines. In higher-dimensional settings, computable lower bounds follow from symmetry, block structure, or sampling of admissible tangent directions. These considerations confirm that the stability constant is not merely existential, but can be estimated effectively from the geometry of the constraint. Related computational aspects of entropy and convex optimization are discussed in [Watrous2018, Wilde2017].
9 Conclusion and outlook
We have established a second-order stability principle for entropy minimization under block-separable convex constraints. The main result, Theorem 4.1, shows that under a confining (fixed-support) hypothesis, the entropy gap controls the squared trace-norm distance to the set of minimizers, with explicit constants determined by the geometry of the constraint.
The analysis reveals a clear geometric mechanism underlying quadratic stability. Although the von Neumann entropy lacks global strong convexity, genuine second-order rigidity emerges locally near constrained minimizers once admissible perturbations are confined to fixed-support strata. The block-separable structure allows the stability problem to decompose into independent marginal and conditional components, whose contributions are governed by the Hessian curvature of Shannon and von Neumann entropy, respectively.
We further show that the quadratic exponent is optimal: even in classical (commutative) settings, no uniform linear or subquadratic stability bound can hold near entropy minimizers. This optimality confirms that the stability mechanism identified here is sharp and intrinsic to the geometry of the constraint.
All results are obtained in finite dimension, where compactness and norm equivalence ensure uniform control of stability constants. Extensions to infinite-dimensional settings or to more general (non block-separable) constraints would require substantially different techniques and are left for future work.
Overall, this work provides a complete and sharp second-order stability theory for entropy minimization under structured convex constraints, with both conceptual and practical implications for optimization and information theory.
Appendix A Domain of validity of quadratic entropy–distance stability
This appendix describes geometric regimes that are explicitly excluded from Theorem 4.1 and explains why a purely quadratic entropy–distance stability estimate cannot hold outside the confining (fixed-support) hypothesis.
Its purpose is not to extend the main result, but to delineate sharply the boundary of its applicability and to show that the failure of quadratic stability in boundary–transversal regimes is structural and unavoidable. This clarification renders the scope of the entropy-based arguments used throughout the paper logically complete and internally consistent.
A.1 Quadratic stability under fixed-support (confining) constraints
Let $\mathcal{C}\subset\mathcal{D}(\mathcal{H})$ be a convex constraint set and let $\sigma\in\mathcal{C}$ be an entropy minimizer. Denote by $P=\operatorname{supp}(\sigma)$ the support projection of $\sigma$.
Lemma A.1 (Quadratic entropy stability in the confining regime).
Assume that there exists a neighbourhood $U$ of $\sigma$ in $\mathcal{C}$ such that
$\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)\qquad\text{for all }\rho\in U.$
Equivalently, admissible perturbations are confined to the fixed-support stratum $\mathcal{D}(P\mathcal{H})$. Then there exist constants $c,C>0$ such that, for all $\rho\in U$,
$c\,\|\rho-\sigma\|_{1}^{2}\;\leq\;S(\rho)-S(\sigma)\;\leq\;C\,\|\rho-\sigma\|_{1}^{2}.$
Proof A.2 (Proof sketch).
On the fixed-support stratum $\mathcal{D}(P\mathcal{H})$, the von Neumann entropy $S$ is a $C^{2}$ functional. Since $\sigma$ is full rank on its support, that is, $P\sigma P\succ 0$ as an operator on $P\mathcal{H}$, the Hessian of f:=Sf:=-S at $\sigma$ is strictly positive on the admissible tangent directions compatible with the constraint. A second-order Taylor expansion therefore yields the stated local quadratic upper and lower bounds in trace norm.
A.2 Failure outside the confining regime
The conclusion of Lemma A.1 relies crucially on the fixed-support (confining) hypothesis. If this assumption fails-specifically, if the entropy minimizer $\sigma$ is rank-deficient and the constraint allows admissible perturbations to explore directions transversal to the support-then the entropy Hessian degenerates at $\sigma$. In such boundary-touching, transversal regimes, the quadratic estimate above is no longer valid.
In that case, the optimal entropy–distance modulus acquires an unavoidable logarithmic correction, and one enters the quadratic–logarithmic rigidity regime analysed in the main body of this work. This distinction explains why purely quadratic entropy-based stability estimates cannot hold uniformly across all symmetry-restricted settings.
A.3 Relation to the main results
Lemma A.1 isolates the geometric regime in which classical quadratic entropy–distance estimates are valid. The main rigidity theorem addresses the complementary geometric regime in which the minimizer remains confined, while boundary–transversal regimes are treated elsewhere.
Together, these results provide a complete geometric picture of entropy stability near constrained minimizers.
Appendix B Standard Entropy Facts
For the reader’s convenience, we collect in this appendix several standard entropy identities and inequalities used in the main text. All results are classical; proofs are included only where brevity permits or where precise hypotheses are required for later arguments.
Throughout, $\mathcal{H}$ denotes a finite-dimensional Hilbert space, and $\mathcal{S}(\mathcal{H})$ the set of positive trace-one operators on $\mathcal{H}$.
B.1 Von Neumann entropy
For $\rho\in\mathcal{S}(\mathcal{H})$, the von Neumann entropy is defined by
$S(\rho):=-\operatorname{Tr}(\rho\log\rho),$
with the convention $0\log 0=0$.
Lemma B.1 (Lower semicontinuity).
The map $\rho\mapsto S(\rho)$ is lower semicontinuous with respect to the trace norm.
Proof B.2.
This follows from Fatou’s lemma applied to the spectral decomposition of $\rho$. See [Wehrl1978, Simon1979].
B.2 Entropy of block-diagonal states
Let $\{P_{i}\}_{i=1}^{r}$ be mutually orthogonal projections with $\sum_{i}P_{i}=I$. A state $\rho$ is block-diagonal with respect to $\{P_{i}\}$ if
$\rho=\sum_{i=1}^{r}P_{i}\rho P_{i}.$
Define $\mu(i):=\operatorname{Tr}(P_{i}\rho)$ and, when $\mu(i)>0$,
$\rho_{i}:=\frac{P_{i}\rho P_{i}}{\mu(i)}.$
Lemma B.3 (Entropy decomposition).
If $\rho$ is block-diagonal with respect to $\{P_{i}\}$, then
$S(\rho)=H(\mu)+\sum_{i=1}^{r}\mu(i)\,S(\rho_{i}),$
where $H(\mu)=-\sum_{i}\mu(i)\log\mu(i)$ is the Shannon entropy.
Proof B.4.
Since the blocks act on orthogonal subspaces, the spectrum of $\rho$ is the union of the spectra of the operators $\mu(i)\rho_{i}$. The claim follows by direct calculation.
B.3 Elementary entropy bounds
We record several standard inequalities relating Shannon entropy to extremal probabilities. All statements below are classical; we include proofs for completeness and to clarify precise constants.
Lemma B.5 (Entropy upper bound).
Let $\mu=(\mu_{1},\dots,\mu_{r})$ be a probability distribution on $r<\infty$ points. Then
$H(\mu)\leq\log r,$
with equality if and only if $\mu_{i}=\frac{1}{r}$ for all $i$.
Proof B.6.
This is classical. The Shannon entropy is strictly concave on the simplex $\Delta_{r}$ and is maximized uniquely at the uniform distribution.
Lemma B.7 (Maximum mass lower bound).
Let $\mu=(\mu_{1},\dots,\mu_{r})$ be a probability distribution. Then
$\max_{i}\mu_{i}\geq e^{-H(\mu)}.$
Proof B.8.
Let $M:=\max_{i}\mu_{i}$. Then
$H(\mu)=-\sum_{i}\mu_{i}\log\mu_{i}\leq-\sum_{i}\mu_{i}\log M=-\log M,$
which implies $M\geq e^{-H(\mu)}$.
Remark B.9.
The bound in Lemma B.7 is sharp, with equality attained for distributions supported on a single point.
Lemma B.10 (Support-size refined bound).
If $\mu$ is supported on at most $r$ points, then
$\max_{i}\mu_{i}\geq\frac{e^{-H(\mu)}}{r}.$
The purely quadratic estimate used here is valid only in the fixed-support (confining) regime; see Appendix A for a precise geometric formulation and proof.
Proof B.11.
By Lemma B.7, $\max_{i}\mu_{i}\geq e^{-H(\mu)}$. Since $\sum_{i}\mu_{i}=1$ and there are at most $r$ nonzero terms,
$1\leq r\cdot\max_{i}\mu_{i},$
which implies $\max_{i}\mu_{i}\geq\frac{1}{r}$. Combining both bounds yields
$\max_{i}\mu_{i}\geq\max\!\left(e^{-H(\mu)},\frac{1}{r}\right)\geq\frac{e^{-H(\mu)}}{r}.$
Remark B.12.
Lemma B.10 is weaker than Lemma B.7 for small entropy, but is convenient when an explicit dependence on support size is needed.
B.4 Trace norm and block structure
Lemma B.13 (Blockwise trace norm).
If $\rho,\sigma$ are block-diagonal with respect to $\{P_{i}\}$, then
$\|\rho-\sigma\|_{1}=\sum_{i=1}^{r}\|P_{i}(\rho-\sigma)P_{i}\|_{1}.$
Proof B.14.
The operators act on orthogonal subspaces, and the trace norm is additive over orthogonal direct sums.
B.5 Pinsker inequality
Lemma B.15 (Pinsker).
For $\rho,\sigma\in\mathcal{S}(\mathcal{H})$ with $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$,
$D(\rho\|\sigma)\geq\tfrac{1}{2}\|\rho-\sigma\|_{1}^{2}.$
Proof B.16.
See [Csiszar1967, Petz2001].
B.6 Use within this work
These results are used explicitly as quantitative entropy inequalities in the proofs of the main theorems:
  • Lemma B.3 underlies the entropy decomposition in Sections 4 and 5.
  • Lemma B.10 is used to bound entropy gaps.
  • Lemma 2.4 yields purity-based lower bounds.
  • Lemma B.13 is used to control trace distances between block-diagonal states.
  • Lemma B.15 is used for relative-entropy comparisons.
No further nontrivial entropy inequalities beyond standard properties are invoked at the level of quantitative estimates. The boundary–transversal regime, which leads to logarithmic corrections, is treated separately and will be discussed elsewhere.
Data Availability
No datasets were generated or analysed during the current study. All results are derived analytically.
Conflict of Interest
The author declares that there is no conflict of interest regarding the publication of this article.
Funding
The author received no external funding.
Code availability
Not applicable.